### v1

# Operating System And Artificial Intelligence: A Systematic Review

Abstract
In the dynamic landscape of technology, the convergence of Artificial Intelligence (AI) and Operating Systems (OS) has emerged as a pivotal arena for innovation. Our exploration focuses on the symbiotic relationship between AI and OS, emphasizing how AI-driven tools enhance OS performance, security, and efficiency, while OS advancements facilitate more sophisticated AI applications. We delve into various AI techniques employed to optimize OS functionalities, including memory management, process scheduling, and intrusion detection. Simultaneously, we analyze the role of OS in providing essential services and infrastructure that enable effective AI application execution, from resource allocation to data processing. The article also addresses challenges and future directions in this domain, emphasizing the imperative of secure and efficient AI integration within OS frameworks. By examining case studies and recent developments, our review provides a comprehensive overview of the current state of AI-OS integration, underscoring its significance in shaping the next generation of computing technologies. Finally, we explore the promising prospects of Intelligent OSes, considering not only how innovative OS architectures will pave the way for groundbreaking opportunities but also how AI will significantly contribute to advancing these next-generation OSs.
1 Introduction
Operating systems (OSes) have long been central to computer systems, efficiently managing hardware resources and providing secure environments for application execution. However, the increasing complexity of modern OSes, the rapid diversification of hardware, and the continuous evolution of Artificial Intelligence (AI) present new opportunities to explore AI’s potential throughout the lifespan of OSes, spanning both development and runtime. As the AI wave continues to surge, AI systems grow increasingly massive and complex, necessitating optimization and efficiency enhancements at the lower layers of the stack, particularly within the operating system.
This review delves into the intricate relationship between AI and OS, examining how AI-driven tools enhance OS performance, security, and efficiency. Conversely, we explore how advancements in OS design facilitate the deployment and optimization of AI applications. By scrutinizing various AI techniques employed to augment OS functionalities—such as memory management, process scheduling, and intrusion detection—this review provides a comprehensive overview of the current state of AI-OS integration. Additionally, we discuss the critical role of the OS in providing essential services and infrastructure necessary for effective AI application operation, from resource allocation to data processing. The article also addresses challenges and future research directions in this domain, emphasizing the imperative for secure and efficient integration of AI capabilities within OS frameworks. Through systematic examination of case studies and recent developments, this review underscores the significance of the AI-OS nexus in propelling the next generation of computing technologies.
In this paper, we embark on a comprehensive exploration of the intersection between AI and OS. Our approach involves collecting and analyzing 108 primary studies in the field, aiming to uncover key insights and trends. The rest of this paper is organized as follows. We provide the research background.Then we propose our research questions, report on the paper selection process, and analyze the distribution of research popularity. Collected papers in the fields of AI4OS and OS4AI are classified and summarized in Sections 4, 5 and 6. Furthermore,we discuss how and what novel OS architectures may create opportunities for AI4OS in section 7 and how LLM can create oppotunies for OS in section 8.
Subsequently, in Section 9, we consolidate these insights by outlining prospective opportunities for future work, particularly emphasizing the synergy between Library OS paradigms and LLM-driven approaches in the pursuit of Intelligent OS. This section underscores the immense potential for synergistic collaborations between these technologies, which could pave the way for highly modular, adaptable, and self-aware operating systems. Finally, we bring this comprehensive exploration to a close in Section 10, synthesizing our findings and offering a compelling conclusion.
2 Background
Operating systems have been providing a crucial layer of abstraction between applications and hardware resources. The design of an OS can significantly impact hardware resource management efficiency and, consequently, the performance of all applications running on it. Additionally, the development of OSes requires extensive engineering efforts from experts. And the complexity of modern OSes poses challenges for developers to effectively improve and optimize performance. The scale and dynamics of computing systems further contribute to this complexity.Moreover, the escalating scale and intricacy of AI systems themselves necessitate optimizations from the OS. As the complexity increases and the need for efficient resource management grows, the integration of AI techniques has emerged as a promising avenue for enhancing OS functionality and performance.
In recent years, researchers have explored various dimensions of AI integration in operating systems, leading to notable advancements.
3 Methodology
We present the research questions, introduce the paper selection process, and statistically analyze the selected papers in this section.
3.1 Research questions
The definition of research questions is the core innovation of a secondary study, as they clearly convey the authors’ perspective on the subject under investigation and the study’s goal.We characterize the purpose, the major research topics, and the scope of this paper as follows.
RQ1: What OS sub-domains are researchers inclined to enhance using AI?
RQ2: How can OS be optimized to improve the efficiency of AI systems?
RQ3: How and which AI techniques are used to improve OS?
RQ4: How and what novel OS architectures create opportunities for AI4OS?
RQ5: How LLMs create opportunities for OS?
3.2 Paper Selection
We identify four AI-related terms and two OS-related terms, then combine these two categories of terms by logical ORs to create nine search strings. After specifying the range the papers were published: 2019–March 2024, we deliver these nine search strings to Google Scholar, to collect related papers in the intersection of AI and OS. The defined search terms are as follows:
AI or Machine Learning or Deep Learning or Large Language Model
And
OS or Operating System
We began with an initial pool of 212 papers related to the intersection of AI and Operating Systems (OS). To narrow down our focus, we applied the following inclusion criteria (IC) and exclusion criteria (EC):
IC1: Peer-Reviewed Research Paper: We considered only peer-reviewed research papers.
IC2: Primary Study: We included primary studies that directly addressed AI-OS integration.
IC3: Relevance to AI and OS: Papers needed to discuss AI techniques enhancing OS or propose novel approaches for improving AI systems using OS methods.
IC4: Publication Date: We restricted our selection to papers published within the past 5 years.
EC1: Avoiding Duplicates: Only the most comprehensive or recent version of each study was included.
EC2: Length Constraint: Papers shorter than 6 pages were excluded.
EC3: Language: We considered papers written in English.
As a result, 108 papers met our criteria and are within the scope of our research. Excluded papers merely mentioned the terms but did not delve into the actual intersection of OS and AI.
Figure 1: Year Distribution of Paper Published
3.3 Paper Analysis
3.3.1 The main venues
Figure 1 depicts the distribution of publication years for research 1. Notably, the number of papers published in the AI4OS and OS4AI domains indicates a growing trend, increased attention in this field. Notably, the number of papers published(including pre-print) in the field of integrating LLM and OS indicate a rapidly growing trend, 17 publications in this field since 2023, when powerful LLM emerged.
Journals and conferences are listed in Tables 1 , the selection criterion is how many times they appear in our paper pool. International Conference on Architectural Support for Programming Languages and Operating Systems(ASPLOS) is the most prevalent conference.
TABLE I: Conferences Distribution
Conference Papers
ASPLOS 22
HPCA 11
USENIX ATC 11
OSDI 5
SOSP 2
HotOS 2
APsys 2
SOSP 2
DAC 1
CCS 1
MobiCom 1
ESEC/FSE 1
ACM TURC 1
ISSTA 1
We first divide all the papers into AI4OS(38), OS4AI(53), LLM AS OS(9) and LLM4OS(8).
4 RQ1: What OS sub-domains are researchers inclined to enhance using AI?
Currently, OSes employ global, static policies based on heuristics. However,AI techniques, adaptable to different application behaviors, hold promise for outperforming these traditional policies. This research question explores and quantifies the diverse applications of AI approaches in enhancing or automating various OS tasks.
4.1 AI for OS Tuning
Growing OS complexity challenges configuration and decision-making, prompting the exploration of AI and ML for enhancements in auto-tuning tasks, including scheduling, energy efficiency, and memory management.
Figure 2: General AI for OS Tuning Workflow
Figure 2 showed the general workflow for AI tools to assist OS.
Scheduling
Scheduling, a core OS function, balances fairness, responsiveness, and throughput by efficiently managing resources.[2][16] Advancements through machine learning can refine ticket distribution, enable adaptability, and adjust scheduling variables precisely.
Springborg et al. [86] introduce Chronus, a Python application that collaborates with the Simple Linux Utility for Resource Management (SLURM) scheduler, prevalent in global supercomputers. Chronus executes comprehensive benchmark tests on HPC clusters, varying parameters like core numbers, processor speeds, and hyperthreading. Throughout, it logs performance metrics and energy use for each setup. This data trains a machine learning model to recognize patterns, aiming to forecast the most energy-efficient configuration for specific jobs and systems. Proven through notable energy savings in benchmarks, this approach holds potential for broader application in HPC systems, promoting eco-friendliness and financial prudence without sacrificing computational power.
Chen et al.[3] presents a ML-based resource-aware load balancer for the Linux kernel with a low-overhead method for collecting training data,an ML model based on a multi-layer perceptron that imitates the Linux’s Completely Fair Scheduler (CFS) load balancer based on the collected training data and an in-kernel implementation of inference on the model. The authors argue that CFS approach maximizes the utilization of processing time but overlooks the contention for lower-level hardware resources.Using eBPF for dynamic tracing, an MLP model replicates CFS decisions, with in-kernel inference for real-time balancing, the model achieves high accuracy and small latency increase, demonstrating effective, low-overhead load management.
Goodarzy et al.[14] also questioned CFS in its ability for proper allocation of CPU, memory, I/O, and network bandwidth. In response to this, the authors propose SmartOS, an operating system that automatically learns what tasks the user deems to be most important at that time. ased on the learned user preferences, SmartOS adjusts the allocation of system resources such as CPU, memory, I/O, and network bandwidth. It prioritizes the resources for the tasks that the user is currently focused on. The authors demonstrate an implementation of such a learning-based OS in Linux and present evaluation results showing that a reinforcement learning-based approach can rapidly learn and adjust system resources to meet user demands.
Storage
Storage systems, along with their associated OS components, are engineered to cater to a broad spectrum of applications and fluctuating workloads. The storage elements within the OS incorporate a range of heuristic algorithms, ensuring optimal performance and adaptability across diverse workloads.
Predictable latency can be very useful for data-center systems serving interactive applications such as messaging and search.Cortez et al.[6] introduce LinnOS, a novel approach to managing SSD performance by incorporating a lightweight neural network within the operating system’s core. This neural network enables real-time inference of SSD performance at an extremely fine-grained level, specifically per-I/O operation, without requiring modifications to the underlying hardware or changes to file systems and applications. LinnOS profiles the latency of a large number of I/O operations submitted to the SSD, leveraging this data to train the neural network. By adopting a ”black-box” perspective towards the device, LinnOS learns and infers the SSD’s behavior to increase predictability, thereby empowering applications to anticipate whether their performance expectations will be met.
Wu et al. [85] introduce LearnWD, a novel approach that harnesses the synergies of machine learning and out-of-place updates to effectively mitigate the write disturbance (WD) problem in NVM. At its core, LearnWD employs clustering algorithms to systematically categorize stale data based on their inherent error proneness. Upon receiving a write request, LearnWD meticulously assesses both the aggressiveness of incoming new data and the error vulnerability of the existing stale data. By thoughtfully leveraging this information, LearnWD strategically orchestrates write operations to proactively minimize the incidence of WD errors.
Wang et al. [87] introduce LearnedFTL,a page-level Flash Translation Layer design that harnesses the power of ML techniques to significantly enhance the random read performance of flash-based SSDs. This method effectively models the non-uniform yet largely linear relationships between logical page numbers and physical page numbers, overcoming the inherent irregularity in flash memory’s address space distribution.LearnedFTL embeds the training of learned indexes within the GC process, allowing for the continuous updating and refinement of the learned models as the address mapping evolves. LearnedFTL also incorporates a bitmap prediction filter acting as a safeguard against potential model inaccuracies, validating the predicted PPNs and ensuring that only correct mappings are used for address translation. This innovative approach uniquely minimizes the occurrence of double reads resulting from address translation during random read accesses.
Memory
Traditional heuristic-based methods struggle due to the intricate hierarchy of memory levels, dynamic workloads, data-dependent behavior, and hardware constraints. These complexities make it difficult to optimize memory management effectively.
Zhang et al.[17] address the challenges of address mapping in 3D-stacking memory, focusing on technologies like High-Bandwidth Memory and Hybrid Memory Cube. They propose Software-Defined Address Mapping, allowing user programs to directly control memory hardware. SDAM enables fine-grained data placement while leveraging chunk-based address mapping management. The system employs machine learning to identify access patterns, resulting in significant speedups compared to fixed address mapping systems.
Lagar-Cavilla et al.[88] introduce a software-defined far memory system that has been successfully deployed across Google’s data centers since 2016. The authors propose a proactive software-defined approach that compresses cold memory pages to create a far memory tier in software. This approach uses the Linux kernel’s zswap mechanism to compress pages and store them in DRAM, effectively creating a far memory tier with low latency. The authors designed an autotuning system that employs a Gaussian Process Bandit to estimate the size of cold memory and the promotion rate under different configurations. This model is used to emulate the control algorithm offline and estimate the impact of different parameter settings. The authors conclude that their software-defined far memory system is effective in saving memory costs without impacting application performance.
Rocha et al. [15] propose PredG, a Machine Learning framework designed to enhance the performance of graph processing. PredG aims to find the ideal thread and data mapping on Non-Uniform Memory Access systems. One of the key features of PredG is its ability to be agnostic to the input graph. It uses the available features of the graphs to train an Artificial Neural Network to perform predictions as new graphs arrive. This is done without any application execution after being trained, which makes it a powerful tool for optimizing graph execution on NUMA machines. This work is part of a broader trend in the research community towards automated graph machine learning, which seeks to discover the best hyper-parameter and neural architecture configuration for different graph tasks/data without manual design.
4.2 AI for OS security
With the widespread application of deep learning in recent years, using deep learning technologies for OS security has emerged, and the effectiveness of threat detection has been dramatically improved.
Qin et al. [35]presents MSNdroid, a novel malware detector designed specifically for Android applications, leverages a combination of native API calls, permissions, system API call features, and a Deep Belief Network. By applying deep learning techniques to native code features, MSNdroid effectively detects Android malware. This approach involves extracting features from a comprehensive dataset comprising malicious applications and benign applications. Notably, MSNDroid achieves an impressive accuracy while maintains an impressively low false-negative rate.
De Wit et al. [39] emphasize the value of incorporating machine learning in malware detection strategies for Android platforms. By leveraging accessible hardware data and sophisticated classification techniques, the study demonstrates the feasibility of identifying malware with a reasonable degree of precision, highlighting the potential of app-specific metrics in enhancing detection rates. This research contributes to the growing body of knowledge on AI-integrated security measures within operating systems, particularly pertinent to the Android ecosystem.
4.3 Findings
In response to Research Question 1, which explores the inclination of researchers to utilize AI in enhancing specific OS sub-domains, our comprehensive analysis of existing literature highlights two prominent areas: auto-tuning and security. These domains have attracted considerable attention due to their potential for transformative improvements through AI integration.
Auto-tuning emerges as a key area where AI is being employed to overcome the shortcomings of conventional heuristic-based methods. These traditional approaches often struggle to cope with the ever-evolving demands of modern computing environments. Researchers are harnessing machine learning to dynamically adjust OS parameters, such as scheduling, energy management, memory allocation, and storage optimization, aiming to boost system performance and resource efficiency. AI’s predictive abilities are particularly crucial in enhancing scheduling mechanisms to strike a balance between fairness, responsiveness, and throughput, effectively managing critical resources, ensuring they can handle the intricate and fluctuating workloads common in today’s computing landscapes.
The security domain has witnessed a surge in the adoption of deep learning technologies to combat malware threats, showcasing substantial advancements in detection efficacy. AI-powered malware detectors have proven superior in terms of accuracy and reduced false negatives compared to traditional machine learning models. Furthermore, AI frameworks are being developed to create adversarial malware, underscoring AI’s dual role in fortifying OS security while also exposing vulnerabilities in AI-based detection systems.
5 RQ2: How can OS be optimized to improve the efficiency of AI systems?
AI accelerators are different from traditional hardware, affecting all aspects of system design, from data-center scale to single-chip scale. They also add high requirement for system architecture, management, and programming[13].
Previous work has shown AI jobs critically demand high-speed I/O and low-latency and high-bandwidth data communication[18][21].Various attempt on hardware has been done to improve ML application performance,for exmaple,using newly NVMe SSD [89], relying on hardware FPGA for the I/O communication control instead of relying on OS-level interrupts that can significantly reduce both total I/O latency and its variance and algorithm level. Accessing hardware through the kernel introduces a performance bottleneck. To mitigate this bottleneck, one effective approach is to bypass the kernel altogether, enabling userspace programs to directly access hardware[20].
Bateni et al.[89] presents NeuOS, a comprehensive system solution aimed at providing latency predictability for executing multi-DNN workloads in autonomous systems while simultaneously managing energy optimization and dynamic accuracy adjustments. It coordinates system- and application-level solutions intelligently to ensure that multiple DNN instances operate efficiently and meet their respective deadlines to guarantee latency predictability. It manages energy consumption by dynamically adjusting parameters such as voltage and frequency scaling to minimize energy usage without compromising the latency predictability or accuracy requirements. Based on specific system constraints, NeuOS adjusts the accuracy level of DNN computations in real-time. This allows for trade-offs between computational precision and resource efficiency, ensuring that the system operates within its given constraints while maintaining an acceptable level of performance.
Wang et al.[21] showed what network for GPU AI remoting,a technique where the execution of GPU APIs is managed remotely through a network on a remote proxy instead of running GPU computations locally on the machine. The study takes a GPU-centric perspective to derive minimum latency and bandwidth requirements and aims to ensure that unmodified AI applications can run on remoting setups using commodity networking hardware without performance degradation. The paper introduces a novel theoretical framework that quantifies the minimum network requirements essential for efficient GPU API remoting. By formalizing the relationship between network latency, bandwidth, and remoting efficiency, this model provides foundational insights.
Serizawa et al. [19]proposed an solution focused on I/O bandwidth.The method aims to improve the reading performance of large training datasets by using high-performance I/O storage devices. The authors discusses the problem of copying datasets between local storage and shared storage and proposes a solution to conceal the time spent on copying by overlapping the copying and reading of training data.
5.1 Findings
Addressing Research Question 2, which delves into the ways OSes can elevate the efficiency of AI systems, our analysis underscores the pivotal function of OSes in handling the distinctive needs of AI tasks.
To optimize AI efficiency through the OS, several strategic approaches emerge as essential. Firstly, there is a need to tailor the OS to cater specifically to the requirements of AI accelerators. This includes facilitating high-speed I/O operations, minimizing latency in communications, and enabling autonomous operation of AI tasks. Secondly, the development of specialized runtime systems and schedulers becomes crucial, ensuring optimal allocation of resources and efficient execution of AI processes. Thirdly,optimizing I/O bandwidth further enhances the performance of AI applications. Collectively, these strategies form a comprehensive framework for improving the efficiency and effectiveness of AI systems through optimized OS integration.
6 RQ3: How and which AI techniques are typically used to improve OS?
To answer RQ3, we review all studies that investigate AI techniques applied in OS in this section. Among 38 AI4OS previous work,those did not clear clarify the AI tools excluded,we list the AI tools that are used more than twice. In general, RF, RNN (LSTM), KNN, RL, MLP,etc. are the most widely-used techniques.
6.1 Tools analysis
RF(DT)
A decision tree is a type of machine learning model used when the relationship between a set of predictor variables and a response variable is non-linear,while random forest is essentially a collection of decision trees. It is quick to fit to a dataset and easy to interpret. Chowdhury et al. [32] used RF as one if the models to accurately detect the attack from the network traffic. To construct the random forest classifier, the authors employed the Random Forest Regressor , which serves as both a regressor and a meta estimator. It accomplishes this by fitting multiple decision trees to different subsets of the dataset. For our specific model, we opted for a forest containing 1000 trees.Ahmed et al. [23] used RF to build a device fitness model,based on the Dataset collected during runtime and statically. De Wit et al.[39] trained a statistical classifier able to recognize malware signatures in any log data collected on a smartphone. The classifier was trained, cross-validated, and tested using the dataset described above and RF classifier had a better performance.Metzge el at.[49] used RF to get a optimal kernel runtime switching slice size. The model is a random forest regressor with 50 decision trees with a depth of two for the GPU model. Ongun et al.[36] used RF to get the probability of a command being malicious,based on labels dataset.
RNN(LSTM)
RNNs are a class of neural networks designed to handle sequential data. They have feedback connections, allowing them to maintain an internal state or memory. Each step in an RNN processes an input and updates its hidden state based on the current input and the previous hidden state.While LSTM ,a type of RNN architecture with a more complex cell structure,were introduced to address the vanishing gradient problem. Motivated by the problem that exploiting 3D-stacking memory’s performance is challenging because bandwidth utilization heavily depends on address mapping in the memory controller, Zhang et al.[17] used a software-defined address mapping, allowing user programs to directly control low-level memory hardware in an intelligent and fine-grained manner.LSTM is used to in a method to get access pattern information to select an address mapping. It identify the major variables that significantly contribute to external memory access and have a substantial impact on memory traffic and data movement.
WordEmbedding
Word embeddings provide a way to achieve this by mapping words to dense vectors in a multi-dimensional space. Fu et al.[34] proposed a AI-based approach to help under-resourced security analysts to find, detect,and localize vulnerabilities. They utilized a word-level Clang tokenizer with a copy mechanism. This tokenizer broke down a C function into a sequence of tokens.Then word embedding was used to generate vector representations for each token in the sequence, capturing the semantic information among the input tokens. And further classifier was , Ongun et al.[36] explore techniques for representing command-line text using word embedding methods.Based on this,they devise ensemble boosting classifiers to differentiate between malicious and non-malicious commands.
KNN
KNN is a supervised machine learning algorithm used for both classification and regression tasks. The fundamental idea behind KNN is simple: neighbors influence each other. If you’re surrounded by similar things, you’re likely similar too. It is widely applicable in pattern recognition, data mining, and intrusion detection. Yang et al[42] introduce a KNN-based machine learning algorithms can accurately predict the Turnaround-time(TaT) of a process. It can effectively reduce the TaT of the process and reduce the number of process context switches.
MLP
An MLP is a type of feedforward neural network used for supervised learning tasks, such as classification and regression. Chen et al.[3] argues that traditional Linux CFS scheduler maximizes the utilization of processing time but overlooks the contention for lower-level hardware resources and try to solve the above problem using an ML-based resource-aware load balancer.They employed supervised imitation learning to replace a portion of its internal logic with an MLP model. This trained MLP model emulates the kernel’s load balancing decisions. MLP is chosen because this current work doesn’t require a very complex model and MLP has a relatively simple implementation compared to the other models. Based on this work,Qiu et al.[78] propose the concept of reconfigurable kernel datapaths that enables kernels to self-optimize dynamically to reduce the cost of kernel. The authors also used MLP ML model that can mimic Linux CFS decisions.
6.2 Findings
In addressing Research Question 3, which explores the methods by which AI techniques are utilized to enhance OS, our analysis reveals a variety of approaches tailored to meet distinct OS requirements. These AI techniques are strategically chosen to address a spectrum of OS challenges, encompassing security enhancements, performance forecasting, resource management, and process scheduling. The widespread adoption of these methods attests to their versatility and underscores the profound influence of AI in augmenting OS functionality and performance. This integration of AI into core OS components not only boosts operational efficiency but also paves the way for more intelligent and adaptable systems, capable of meeting the evolving demands of modern computing environments.
7 RQ4: How and what novel OS architectures create opportunities for AI4OS?
Comprehensive research has revealed that utilizing the standard kernel pathway for hardware interaction results in notable efficiency losses, rendering conventional approaches less than ideal for the current AI-driven environment. In response, kernel bypass tactics are gaining prominence, aiming to optimize hardware utilization and enhance real-time capabilities specifically for AI tasks. Nonetheless, the absence of an OS means these strategies fall short in providing the necessary tailoring to exploit AI applications’ full potential.
In essence, the interplay between modern application design and OS innovation fosters a fertile ground for the conception of systems that not only meet but also anticipate the evolving needs of AI-driven environments. This confluence of advancements signals a pivotal moment in the trajectory of computing, where the synergy between AI and OS architectures could redefine the boundaries of what is achievable in high-performance computing.
7.1 Kernel-bypass OS Structure for AI
Prior research has successfully demonstrated the potential of leveraging hardware acceleration for machine learning within the kernel space, showcasing the feasibility of such an approach[4]. This study introducES an API remoting system, which facilitates access to specialized accelerator interfaces for kernel space applications. Moreover, it simplifies the integration by offering high-level APIs directly to the kernel space, eliminating the necessity for kernel-specific adaptations of complex libraries. The API remoting mechanism transmits commands between kernel and user space. This innovative design not only enhances the performance of AI applications within the kernel but also provides a compelling perspective on the capabilities of modern operating systems.
Raza et al. [53] propose integrating unikernel optimizations into Linux, known for creating secure, compact OS images for single applications. Unikernel Linux (UKL) reduces the number of executed instructions and improves instructions-per-cycle efficiency. Tail latency tests on Memcached—a multi-threaded key-value store—show that UKL achieves a significant performance improvement. It introduces a configuration option that allows a single, optimized process to link directly with the kernel, bypassing the traditional system call overhead, and significantly cuts latency for system call payloads, showcasing the benefits of unikernel-inspired enhancements in refining Linux’s performance.
Cadden et al.[59] introduces Serverless Execution via Unikernel SnapShots (SEUSS), a method that leverages unikernel snapshots for rapid deployment and high-density caching of serverless functions. The authors describe the use of unikernel contexts, which consist of a high-level language interpreter configured to import and execute function code, providing isolation and security. This minimalistic approach leads to a reduced memory footprint and faster startup times compared to traditional operating systems like Linux, which is beneficial for serverless environments where rapid function instantiation is crucial. By using unikernels, the SEUSS system can cache a large number of function instances in memory due to the reduced memory footprint.
7.2 Library OS
One significant obstacle that OSs face in effectively leveraging the potential of AI lies in the absence of a comprehensive, universally applicable strategy for adapting AI technologies to the wide array of heterogeneous devices in use.
The Demikernel project[52] unveils an OS architecture optimized for datacenter systems with microsecond-scale requirements, emphasizing low-latency I/O. Adopting a LibOS strategy, it side-steps the traditional kernel in I/O paths, significantly enhancing performance. This design supports kernel-bypass devices, allowing seamless application operation with negligible overhead at the nanosecond scale. For I/O efficiency, Demikernel applies zero-copy techniques for large buffers, optimizes resource use, and maintains system stability through periodic LibOS interaction. Security is bolstered with controlled data placement and potential advanced memory integrity measures. LibOS components employ hardware acceleration to offload critical tasks, minimizing latency and maximizing throughput. Kernel bypass in I/O paths reduces overhead from context switching, system calls, and memory duplication.
7.3 Other Stuctures
Skiadopoulos et al. [60] present DBOS as a superior alternative to conventional cluster OS components, offering comparable functionality but enhanced analytics and reduced code complexity. DBOS matches current systems in scheduling, file management, and inter-process communication, yet excels in analytics and simplifies code through database query-based OS services. It efficiently implements low-latency transactions and ensures high availability. DBOS’s integrated DBMS approach is especially advantageous for ML, delivering a cohesive platform for efficient resource management and analytics in large-scale distributed environments, adeptly managing parallel computation and workload dynamics across various hardware.
Shan et al.[61] developed LegoOS—a splitkernel architecture for hardware disaggregation, decentralizing OS functions for scalable, distributed management. It fundamentally breaks down traditional OS functionalities into loosely-coupled monitors that each run on and manage a distinct hardware component. The splitkernel model distributes responsibilities such as scheduling, memory management, and I/O operations across these monitors, effectively creating a distributed set of hardware components that can be independently managed and scaled. The splitkernel model in LegoOS allows for independent scaling of compute and memory resources, essential for ML scenarios where large datasets and complex models necessitate specialized hardware accelerators.
7.4 Findings
Addressing Research Question 4, our analysis highlights the transformative role of novel OS architectures in advancing AI4OS. Pioneering designs like LibOS and unikernels have sparked renewed excitement among researchers, demonstrating their potential to outshine traditional OSes. These cutting-edge architectures enhance the symbiosis between software and hardware, reducing latency to boost AI application responsiveness and real-time processing capabilities. Their streamlined construction optimizes resource allocation, ensuring efficient management of computational assets. Beyond performance, these architectures establish a secure foundation for AI deployment. By simplifying the environment, they minimize security risks, vital for AI applications managing sensitive data. DBOS and LegoOS exemplify this by simplifying distributed AI setup and maintenance, freeing experts to focus on algorithmic innovation. They also adapt nimbly to the variable demands and hardware diversity of modern ML, enhancing deployment efficiency.
In essence, these advanced OS architectures elevate performance and security, fostering an innovative landscape for AI4OS. By catering to AI application requirements, they promise to rewrite the rules of computing, blending AI and OS synergies for smarter, more efficient, and secure computing futures.
8 RQ5: How LLMs create opportunities for OS?
The integration of LLMs into OSes presents a significant opportunity to enhance the user experience and overall system functionality[67][68]. LLMs, with their advanced natural language processing capabilities,[62][63][64] can transform the way users interact with and manage their computing environments.[76] We simply divide these works into two big categories: LLM AS OS and LLM4OS.
8.1 LLM AS OS
”LLM AS OS” refers to the integration of LLMs into the core of an OS, effectively serving as the ”brain” of the system,make OS capable of understanding and responding to natural language commands, thereby enabling more intuitive and flexible human-computer interaction[70][69].
Kamath et al.’s LLaMaS[54] addresses the complexities of diverse computing environments by leveraging LLMs to ease OS challenges in hardware adaptation and resource management. It adapts to new devices by interpreting plain text specs, recommending optimized OS tactics. The LLM frontend deciphers system characteristics, converting them into actionable embeddings. The backend uses these for on-the-fly OS decisions, like memory tier data movement. An experiment with ChatGPT demonstrated its skill in adjusting memory allocation for CPU and GPU tasks based on usage. Aimed at reducing administrative and research costs, LLaMaS autonomously aligns with hardware changes via LLMs’ innate zero-shot learning, eliminating the need for manual adjustments or detailed device-specific coding.
AIOS-Agent ecosystem[55][56], where LLMs act essentially as an ”operating system with soul”.LLMs are embedded in the OS kernel for intelligent decision-making and resource allocation. Its context window acts as memory, managing relevant data, while external storage serves as a file system with enhanced retrieval capabilities. Hardware and software are treated as peripherals and libraries, enabling agent-environment interaction. Natural language becomes the primary programming interface, democratizing software development. This ecosystem supports single and multi-agent applications for executing a broad range of tasks.
MemGPT[57] presents a solution to overcome the fixed-length context window limitation in LLMs, a hurdle for tasks demanding deep analysis of lengthy dialogues or extensive texts. Drawing parallels with traditional OS hierarchical memory systems, MemGPT introduces virtual context management, echoing virtual memory principles. It enables LLMs to access information beyond immediate capacity, mirroring OS memory management through strategic ’paging’. This system reacts to events like user messages, system alerts, and timers, appending pertinent data to the primary context buffer before processing. Additionally, it supports function chaining for uninterrupted, sequential task execution, accommodating intricate operations and long-term planning within the LLM’s context limitations.
8.2 LLM4OS
SecRepair[58], designed to address the challenge of identifying and repairing code vulnerabilities in software development. This system is powered by a LLM and incorporates reinforcement learning and semantic rewards to enhance its capabilities.To support the training process and prepare a robust model for vulnerability analysis, the authors have compiled and released a comprehensive instruction-based vulnerability dataset.They also propose a reinforcement learning technique with a semantic reward mechanism to generate concise and appropriate code commit comments. This technique is inspired by how humans fix code issues and provides developers with a clear understanding of the vulnerability and its resolution.
Rahman et al. [71] introduce ChronoCTI, an automated pipeline for mining temporal dynamics between attacker actions from text-based accounts of cyber incidents. The focus is on pinpointing repetitive action sequences, termed temporal attack patterns, crucial for preemptive defense strategies by security professionals against impending cyber threats. ChronoCTI is anchored by a curated dataset linking sentences to adversary maneuvers, temporal linkages across 94 attack narratives, and large language models refined on cybersecurity topics. The research delineates a structured approach, harnessing cutting-edge language models, NLP methods, and ML techniques to decipher the temporal structure within attack stories.
8.3 Findings
Addressing Research Question 5, our analysis highlights the incorporation of LLMs within operating systems heralds a transformative era, reshaping system dynamics and enhancing user engagement. They can enrich user interfaces by facilitating natural language processing for voice commands and text inputs, leading to more intuitive and personalized user experiences. In the realm of development, LLMs can automate code generation and optimization, streamlining coding processes and reducing errors. Additionally, they contribute to dynamic system management by analyzing usage patterns to optimize resource allocation, boosting system responsiveness. LLMs also play a pivotal role in intelligent troubleshooting, diagnosing system issues more accurately and recommending preventive measures to minimize downtime. Furthermore, they bolster OS security by detecting anomalies indicative of malicious activities and assisting in secure configuration management. Lastly, LLMs act as intelligent aids for developers, offering insights into function queries, suggesting best practices, and even generating documentation, thus accelerating the development process and enhancing code quality. Through continuous learning, LLMs ensure that OS performance and efficiency improve over time, adapting to user needs and operational contexts.
9 Opportunities for Future Work
9.1 Optimized OS for AI
The escalating sophistication of AI workloads, characterized by high-velocity I/O, minimal latency, and bandwidth-intensive requirements, is catalyzing an urgent need for AI-specialized OS optimization. As AI permeates diverse sectors, the imperative for a finely calibrated infrastructure escalates, poised to support the burgeoning demands of next-generation applications. The emergence of AI-dedicated hardware has dramatically reshaped AI computation landscapes. These technological marvels, while delivering unrivaled performance, necessitate software ecosystems meticulously attuned to their singular attributes.
In this multifaceted hardware ecosystem [20][23], the need for an OS that can seamlessly manage and orchestrate AI tasks across a spectrum of devices is more pressing than ever. AI has transcended beyond the realm of traditional neural networks, with innovative models leading the change in natural language processing and graph neural networks adeptly handling intricate relational data. These advanced applications are calling for specialized optimizations that a one-size-fits-all approach cannot provide. Juggling performance, energy conservation, and scalability constitutes a multifaceted challenge, demanding an OS equipped for dynamic adjustment to cater to the singular requisites of every workload.
Portatble OS for AI
In industrial settings, IoT devices generate massive amounts of data. Deploying deep learning models directly on edge devices (such as sensors, gateways, or edge servers) in equally important as relying solely on cloud-based processing. In a certain sense, deploying on edge devices can be a more challenging scenario because the computational power and bandwidth, of edge devices are more constrained[77][12][1]. Due to the diversity of edge devices,current AI platforms’ lack of portability across different edge platforms hinders widespread adoption ,and bypassing the kernel and directly accessing hardware while adapting to various hardware configurations can be extremely challenging. Therefore, an operating system that can accommodate multiple hardware types and is specifically tailored for AI applications is a potential solution[7][8].
Fast deployment and standard API interface
Simplified deployment means that even non-experts can utilize AI applications effectively. Developers, data scientists, and industrial engineers can quickly integrate the framework into their existing edge devices without extensive configuration or manual setup. And a standard API interface ensures that software components can communicate seamlessly with each other.
Minimized OS consumption
Resource-constrained environments, such as edge devices or embedded systems, often operate with tight constraints on memory, processing power, and energy while in scenarios like cloud computing or data centers, minimizing resource consumption directly impacts operational expenses. As OS and applications share the resources, the OS consumption determines the actual operational performance of the algorithmic model. So there is a need to design an efficient OS that can achieve small memory and fast training and inference. There are several promising approaches to achieve this. Apart from bypassing the kernel, unikernel OSes are designed to be fast and lightweight[11][32].
Privacy and security
The problem of privacy preservation in the context of machine learning is quite different from that in traditional data privacy protection, as machine learning can act as both friend and foe. Increasingly, edge devices are equipped with AI applications. This trend comes with privacy risks as models can leak information about their training data [24]. OS need to find ways to provide safeguards for ensuring the confidentiality and integrity of its data and programs.
9.2 Intelligent OS
9.2.1 Potential for Intelligent LibOS
Figure 3: Monolithic application to Microservice: ① Microservices enable independent, scalable deployments ② Microservices provide fine-grained security controls ③ Microservice enables fine-grained flexible upgrading and replacement of individual microservice ④ Microservice enables service intelligent orchestration ⑤ Microservice allows agile and independent development for each service
Figure 4: Monolithic OS to LibOS:①LibOS spans across multiple hosts.②LibOS offers flexible, granular security and kernel-bypass capabilities.③Only LibOS supports granular updates to individual components.④LibOS enables orchestrating components intelligently for adaptable system contexts.⑤LibOS facilitates agile, independent development of single components.
A recently emerging trend of Internet-based software systems is “resource adaptive,” i.e., software systems should be robust and intelligent enough to the changes of heterogeneous resources[9], both physical and logical, provided by their running environment[26]. The key principle of such an OS is based on resource disaggregation, resource provisioning as a service, and learning-based resource scheduling and allocation. Meanwhile, the OS should leverage advanced machine/deep learning techniques to derive configurations and policies and automatically learn to tune itself and schedule resources. One general and efficient Spark tuning framework is proposed and applied it to Tencent’s data platform[5].
LibOS have risen to prominence as highly attractive options in scenarios that prioritize agility and robust security, thanks to their inherent minimalist and modular architectures. Figure 3 illustrates the progression from monolithic applications to microservices, and figure 4 the contrast between a monolithic OS and LibOS.
Microservices excels notably in granular security, isolated updates, and performance optimization through dynamic service management. AI technologies have proven adept at governing microservices, enabling intelligent service orchestration and customization, as highlighted in research [65] and [66], setting the stage for advanced system control and optimization[81].LibOS similarly exhibit gains in agility, composability, enhanced isolation, and facilitated deployment across heterogeneous hardware. LibOS presents a more auspicious ground for intelligent OS. There are at least 5 ways where LibOS will benefit to enable a Intelligent OS:
Isolation and security
FlexOS[82] presents a modular LibOS framework, comprised of discrete elements that can be segregated using a spectrum of hardware safeguards, complemented by adaptable data sharing protocols and software fortification methods. FlexOS adopts micro-libraries as fundamental units for segmentation. Integration or isolation of these components is dynamically managed, guided by safety and performance requirements.
Drawing from the microservices paradigm’s strengths in isolation and precise access governance, we posit that LibOS offers a strategic avenue to uphold data confidentiality during processing and safeguard overall system integrity. In the AI domain, where security and privacy are paramount[37]. LibOS’s robust isolation features make them ideally suited for secure operation in environments with concurrent tasks of varying privacy sensitivity. The innate simplicity and modularity of LibOS create an ecosystem that effectively quarantines critical workloads, bolstering confidence and aligning with rigorous privacy statutes.
Heterogeneous deployment
In the IoT epoch, a surge in varied endpoint devices underscores the imperative for AI integration. Yet, conventional monolithic OS kernels struggle, weighed down by excessive resource demands and challenges in hardware acclimatization, hindering their agility in today’s rapidly transforming milieu. In alignment, LibOS surface as a beacon of hope for versatile, cross-platform deployment amid the AI revolution. The Demikernel project[52] substantiates LibOS’s flexibility, enabling applications to operate fluidly across disparate devices sans customization. In LibOS’s schema, hardware interface components are architected as pluggable modules, amplifying interoperability and malleability.The intrinsic modularity and adaptability simplify AI assimilation into a spectrum of devices, nurturing a more dynamic and reactive IoT framework.
Intelligent OS orchestration
A significant advancement brought about by the microservices paradigm is its capacity for orchestration for components,exemplified by tools like ISTIO. This capability serves as an inspiration for the realization that similar traffic control principles can be harnessed within the realm of LibOS to achieve intelligent resource management. In the context of demanding tasks such as machine learning training that consume vast amounts of resources, as well as in resource-constrained edge devices, meticulous oversight over resource allocation is paramount. This empowers dynamic, autonomous resource allocation, optimizing performance and efficiency across varied workloads and settings, perfectly suiting complex, evolving AI applications and infrastructures.
Intelligent OS library replacement
The modular LibOS architecture promotes autonomous component operation through precise interfaces, ensuring loose coupling and stable system evolution akin to microservices. Components can be independently upgraded, patched, or optimized, minimizing disruption and extensive testing. Upon detecting issues like performance drops, an orchestration layer automates replacements, preserving system robustness, security, and adaptability without compromising user experience or structural coherence.
Intelligent OS development
Microservices architecture heralds a seismic shift in software development paradigms, dismantling centralized structures in favor of agile, decentralized teams, each specializing in a unique microservice[83]. This transition boosts efficiency and reliability, as focused teams oversee discrete services, revolutionizing software engineering approaches. Contrastingly, monolithic systems’ vast scale and tangled dependencies escalate code generation complexities. Automating code in large, interconnected systems demands meticulous integration within a dense module network[84]. LibOS, with their compact, independent nature, simplify code generation, offering fertile ground for efficient automation and AI-driven development.
9.2.2 Multi-agent LLM for OS Development
Previous research has demonstrated the efficacy of employing AI in optimizing, safeguarding, and evaluating system software including operating systems[10][22][28]. Our study reveals a notable dearth of efforts dedicated to leveraging LLMs in the context of operating systems. Nonetheless, we identify at least one distinct domains where LLMs can potentially revolutionize the functioning and capabilities of OSes as Multi-agent LLM for OS Development.
Figure 5: One example for Multi-agent LLM for OS Development
Recent breakthroughs in LLM-based multi-agent systems, building on the success of single LLMs as unified decision-makers, have unlocked new horizons in complex problem-solving and environment simulation. This progress signals the potential for multi-agent LLMs to revolutionize OS development. Studies like [79] and [80] showcase multi-agent LLM models’ efficacy in software development, from analysis to debugging.
In the realm of OS development, the integration of multi-agent systems fosters a specialized, collaborative, and flexible environment. A constellation of agents, each dedicated to niches such as kernel architecture, driver engineering, or cybersecurity, collectively enhances project productivity[72][73][74][75]. these agents facilitate knowledge exchange and consensus-building akin to human collaboration. They concurrently tackle OS components, recalibrate in response to evolving project dynamics, and oversee integration, testing, and debugging phases, proactively addressing potential issues.Inherently adaptable, the system reassigns agent responsibilities or integrates fresh entities to align with technological advancements, market dynamics, or emerging security vulnerabilities, ensuring continuous OS refinement. Certain agents conduct simulations to assess OS performance under varied scenarios, supplying empirical insights to inform judicious decision-making, preserving the OS’s technological relevance and competitiveness.
9.2.3 LLM AS OS
Future research in the domain of LLMs integrated as OS will focus on following areas.
Platform-Agnostic Adaptation
LLMs’ inherent understanding of natural language empowers them to interpret a spectrum of system artifacts, from logs to hardware specs, irrespective of software platforms or hardware architectures. This capability facilitates dynamic OS parameter adjustment and context-aware optimization strategies, ensuring the system stays in sync with hardware advancements and performs optimally across diverse ecosystems.
Intelligent Resource Allocation
Leveraging pattern recognition and predictive analytics, LLMs can make informed real-time resource management decisions, preemptively addressing resource demands and workload fluctuations. This proactive approach optimizes system performance, reduces latency, and maximizes resource utilization.
Natural Language Interface
Integrating LLMs into the OS interface enables a conversational interaction model, replacing traditional interfaces with intuitive text-based requests. This simplifies user engagement and broadens accessibility, enhancing user experience and lowering the learning curve.
Customized User Experience
LLMs learn from user behavior and preferences, dynamically adjusting the OS interface and functionality to cater to individual needs. This personalization fosters a more engaging and productive user interface, tailored to unique usage patterns.
Enhanced Security Measures
LLMs contribute to OS security by analyzing data for threats, enforcing policies, and managing access. They can predict security incidents, provide real-time code analysis, and offer personalized security guidance, fortifying the system against vulnerabilities.
10 Conclusion
The synthesis of AI and OS is reshaping the landscape of contemporary computing, as evidenced by our meticulous review of 108 primary studies. AI’s integration into OS frameworks has catalyzed significant improvements in key areas like memory oversight, task coordination, and security vigilance. By leveraging intelligent decision-making beyond traditional rule-bound paradigms, AI bolsters resource efficiency and system agility.
In tandem, advanced OS designs are morphing to better serve AI applications, especially as data-intensive workloads surge across various computing platforms. The OS’s evolving role in adeptly handling specialized hardware and optimizing data flow underscores its importance in sustaining AI software performance.
The integration of LLMs into OS interfaces, as highlighted in our analysis of 17 studies, marks a pivotal advancement in user experience. Conversational interfaces enable seamless human-computer interaction, simplifying device management and personalizing user assistance through natural language processing.
Despite the optimistic outlook, integrating AI into OS presents critical challenges. Safeguarding against security breaches and privacy infringements while maintaining efficient AI operations necessitates rigorous design considerations. Moreover, the demand for streamlined AI algorithms that can operate within constrained environments remains a high-priority research area.
Future research agendas should concentrate on the domain of intelligent operating systems, capitalizing on the transformative potential they offer. Central to this endeavor is the exploration of LibOS. With its modular structure, LibOS facilitates the seamless incorporation of AI enhancements, supports granular control, fosters innovation, and upholds stringent security standards. LibOS stands as a robust foundation for the evolution of intelligent computing. Thus, advancing LibOS technologies should be a focal point for researchers aiming to realize the full scope of intelligent OS capabilities and drive the computing industry toward greater adaptability, efficiency, and resilience.
Acknowledgments
This work was supported in part by the Major Program of National Natural Science Foundation of Zhejiang(LD24F020014), and in part by the Zhejiang Pioneer (Jianbing) Project (2024C01032), and in part by the Key R&D Program of Ningbo(2023Z235), and in part by the Ningbo Yongjiang Talent Programme(2023A-198-G).

### v2

@@
-Operating System And Artificial Intelligence: A Systematic Review
+Integrating Artificial Intelligence into Operating Systems: A Comprehensive Survey on Techniques, Applications, and Future Directions
 Abstract
-In the dynamic landscape of technology, the convergence of Artificial Intelligence (AI) and Operating Systems (OS) has emerged as a pivotal arena for innovation.
-Our exploration focuses on the symbiotic relationship between AI and OS, emphasizing how AI-driven tools enhance OS performance, security, and efficiency, while OS advancements facilitate more sophisticated AI applications.
-We delve into various AI techniques employed to optimize OS functionalities, including memory management, process scheduling, and intrusion detection.
-Simultaneously, we analyze the role of OS in providing essential services and infrastructure that enable effective AI application execution, from resource allocation to data processing.
-The article also addresses challenges and future directions in this domain, emphasizing the imperative of secure and efficient AI integration within OS frameworks.
-By examining case studies and recent developments, our review provides a comprehensive overview of the current state of AI-OS integration, underscoring its significance in shaping the next generation of computing technologies.
-Finally, we explore the promising prospects of Intelligent OSes, considering not only how innovative OS architectures will pave the way for groundbreaking opportunities but also how AI will significantly contribute to advancing these next-generation OSs.
+In the era of the Internet of Everything, operating systems (OSs) face unprecedented challenges posed by an evolving application landscape and increasingly heterogeneous hardware ecosystems.
+This shift toward increasingly dynamic and unpredictable operational contexts presents significant challenges for both OS developers and users.
+Against this backdrop, the fusion of Artificial Intelligence (AI) with Operating Systems emerges as a critical frontier for innovation.
+This survey delves into the intricate interplay between AI and OSs, illustrating how existing OS mechanisms combined with AI significantly elevate the performance, security, and efficiency of modern operating systems.
+We investigate a range of AI methodologies applied to optimize core OS functionalities and clarify the correlation to related studies.
+Our analysis touches on the existing hurdles and prospective avenues in this interdisciplinary domain, underscoring the imperative for robust and seamless integration of AI capabilities into OS architectures.
+Through an examination of illustrative case studies and cutting-edge developments, we offer a thorough review of the current status of AI-OS integration, accentuating its pivotal role in steering the evolution of advanced computing paradigms.
+We also envision the promising prospects of Intelligent Operating Systems, debating how groundbreaking OS designs will usher in novel possibilities and highlight the central role that AI will assume in propelling these next-generation systems forward.
+This forward-thinking outlook illuminates the profound influence of AI on the foundational elements of computing, heralding the advent of a new epoch characterized by intelligent, self-adapting, and highly adaptive software ecosystems.
 1 Introduction
-Operating systems (OSes) have long been central to computer systems, efficiently managing hardware resources and providing secure environments for application execution.
-However, the increasing complexity of modern OSes, the rapid diversification of hardware, and the continuous evolution of Artificial Intelligence (AI) present new opportunities to explore AI’s potential throughout the lifespan of OSes, spanning both development and runtime.
-As the AI wave continues to surge, AI systems grow increasingly massive and complex, necessitating optimization and efficiency enhancements at the lower layers of the stack, particularly within the operating system.
-This review delves into the intricate relationship between AI and OS, examining how AI-driven tools enhance OS performance, security, and efficiency.
-Conversely, we explore how advancements in OS design facilitate the deployment and optimization of AI applications.
-By scrutinizing various AI techniques employed to augment OS functionalities—such as memory management, process scheduling, and intrusion detection—this review provides a comprehensive overview of the current state of AI-OS integration.
-Additionally, we discuss the critical role of the OS in providing essential services and infrastructure necessary for effective AI application operation, from resource allocation to data processing.
-The article also addresses challenges and future research directions in this domain, emphasizing the imperative for secure and efficient integration of AI capabilities within OS frameworks.
-Through systematic examination of case studies and recent developments, this review underscores the significance of the AI-OS nexus in propelling the next generation of computing technologies.
-In this paper, we embark on a comprehensive exploration of the intersection between AI and OS.
-Our approach involves collecting and analyzing 108 primary studies in the field, aiming to uncover key insights and trends.
-The rest of this paper is organized as follows.
-We provide the research background.Then we propose our research questions, report on the paper selection process, and analyze the distribution of research popularity.
-Collected papers in the fields of AI4OS and OS4AI are classified and summarized in Sections 4, 5 and 6.
-Furthermore,we discuss how and what novel OS architectures may create opportunities for AI4OS in section 7 and how LLM can create oppotunies for OS in section 8.
-Subsequently, in Section 9, we consolidate these insights by outlining prospective opportunities for future work, particularly emphasizing the synergy between Library OS paradigms and LLM-driven approaches in the pursuit of Intelligent OS.
-This section underscores the immense potential for synergistic collaborations between these technologies, which could pave the way for highly modular, adaptable, and self-aware operating systems.
-Finally, we bring this comprehensive exploration to a close in Section 10, synthesizing our findings and offering a compelling conclusion.
-2 Background
-Operating systems have been providing a crucial layer of abstraction between applications and hardware resources.
-The design of an OS can significantly impact hardware resource management efficiency and, consequently, the performance of all applications running on it.
-Additionally, the development of OSes requires extensive engineering efforts from experts.
-And the complexity of modern OSes poses challenges for developers to effectively improve and optimize performance.
-The scale and dynamics of computing systems further contribute to this complexity.Moreover, the escalating scale and intricacy of AI systems themselves necessitate optimizations from the OS.
-As the complexity increases and the need for efficient resource management grows, the integration of AI techniques has emerged as a promising avenue for enhancing OS functionality and performance.
-In recent years, researchers have explored various dimensions of AI integration in operating systems, leading to notable advancements.
-3 Methodology
-We present the research questions, introduce the paper selection process, and statistically analyze the selected papers in this section.
-3.1 Research questions
-The definition of research questions is the core innovation of a secondary study, as they clearly convey the authors’ perspective on the subject under investigation and the study’s goal.We characterize the purpose, the major research topics, and the scope of this paper as follows.
-RQ1: What OS sub-domains are researchers inclined to enhance using AI?
-RQ2: How can OS be optimized to improve the efficiency of AI systems?
-RQ3: How and which AI techniques are used to improve OS?
-RQ4: How and what novel OS architectures create opportunities for AI4OS?
-RQ5: How LLMs create opportunities for OS?
+The evolution of operating systems (OS) has been shaped by the continuous interplay between technological advancements and the changing demands of computing environments.
+From the early days of manually configured systems to today’s sophisticated platforms, OS development has been driven by the need to efficiently manage resources, enhance usability, and adapt to emerging paradigms.
+Key milestones, such as the introduction of Unix and the rise of graphical user interfaces, laid the foundation for modern operating systems, enabling multitasking, modularity, and accessibility.
+Over the decades, the diversification of computing devices, from desktops and mobile devices to embedded systems, has further propelled innovations in OS design to address the unique requirements of specialized hardware and applications.
+In recent years, the integration of artificial intelligence (AI) into operating systems has emerged as a transformative trend, unlocking new opportunities for intelligent resource management, adaptive performance, and enhanced user experiences.
+AI encompasses a range of technologies, including expert systems, machine learning (ML), and large language models (LLMs), each contributing distinct capabilities to OS functionality.
+By embedding AI technologies, operating systems can evolve from static, rule-based platforms into dynamic, intelligent systems capable of optimizing operations, predicting user needs, and addressing the growing complexity of modern computing environments.
+The motivation for integrating AI into OS lies in its potential to address critical challenges in areas such as resource allocation, security, and system adaptability.
+For example, ML algorithms enable OSs to learn from data and adjust configurations dynamically, while LLMs enhance user interactions through natural language processing, enabling intuitive and conversational interfaces.
+Additionally, the adoption of AI-driven techniques aligns with emerging trends in computing, such as the Internet of Things (IoT), edge computing, and cloud systems, where managing large-scale, interconnected environments requires intelligent and scalable solutions.
+This survey explores the intersection of AI and OS, providing a comprehensive review of recent advancements, challenges, and future opportunities.
+We categorize the integration of AI technologies into two key aspects of OS development: design and operational optimization.
+Our analysis is based on an extensive review of primary studies, highlighting the impact of AI on OS performance, efficiency, and adaptability.
+The remainder of this paper is organized as follows: Section 2 outlines the scope and objectives of the survey.
+Sections 3 and 4 discuss AI-driven methodologies for OS development and operational enhancements.
+Section 5 examines the integration of LLMs into OS architectures, and section 6 discusses the clarification of different stages of AI-assisted OSs.
+Section 7 identifies emerging trends and future directions, while Section 8 concludes the paper, summarizing the key insights and contributions.
+Here are the highlights of this paper:
+  • We identify and discuss the application of ML tools across various subdomains of operating systems.
+  • We provide detailed insights into a variety of ML algorithms used in different areas of operating system research, which includes understanding the specific methodologies employed to achieve performance improvements and security enhancements.
+  • We explore how LLMs can be utilized within the context of operating systems, suggesting new research avenues and potential improvements that can be realized through their integration.
+  • We identify novel operating system architectures that provide opportunities for the development of intelligent operating systems.
+This includes examining designs that facilitate better AI integration and enable advanced features such as dynamic resource allocation and improved security protocols.
+  • We define and outline the different stages of AI integration within operating systems, providing a roadmap for ongoing and future research efforts aimed at enhancing OS functionality through AI.
+2 Related Search
+Previous research has made significant strides in exploring the integration of AI with OS (?).
+For instance, (?) examines the application of quantum computing in cybersecurity, highlighting how AI and qubit-based operating systems can enhance security measures.
+This work points to the potential of quantum-enhanced systems to effectively address complex security challenges.
+Additionally, (?) investigates the use of object-oriented programming and AI methodologies to structure and optimize operating systems, advocating for more compact, customized systems that meet the diverse and evolving demands of various computing environments.
+This study underscores the importance of adaptability and efficiency in modern OS design.
+In (?), the focus is on the role of AI and ML in OS management, presenting a framework for their integration aimed at enhancing resource allocation, security, performance, and system adaptability.
+This research emphasizes the need for intelligent systems capable of self-optimization and dynamic responses to changing conditions.
+Furthermore, (?) assesses the impact of various AI techniques on key aspects of distributed operating systems, including resource management, fault tolerance, and security.
+(?) focuses on the improvements that ML can bring to OS resource management, while (?), (?), and (?) address security aspects of operating systems.
+Additionally, (?) explores the application of AI methods in software configurations.
+While these works have provided valuable insights and advancements, they often concentrate on specific types of operating systems or may now be considered somewhat outdated, particularly given the recent emergence and rapid development of large language models (?).
+As we move forward, it will be crucial to explore how LLMs and other cutting-edge AI technologies can be integrated into operating systems to drive the next wave of innovation in computing.
+3 Research Method
 3.1 Paper Selection
-We identify four AI-related terms and two OS-related terms, then combine these two categories of terms by logical ORs to create nine search strings.
-After specifying the range the papers were published: 2019–March 2024, we deliver these nine search strings to Google Scholar, to collect related papers in the intersection of AI and OS.
-The defined search terms are as follows:
-AI or Machine Learning or Deep Learning or Large Language Model
+3.1.1 Search string design
+We identify five AI-related terms and two OS-related terms, then combine these two categories of terms by logical ORs to create ten search strings.
+AI or Machine Learningor Deep Learning or Large Language Model or LLM
 And
 OS or Operating System
-We began with an initial pool of 212 papers related to the intersection of AI and Operating Systems (OS).
-To narrow down our focus, we applied the following inclusion criteria (IC) and exclusion criteria (EC):
-IC1: Peer-Reviewed Research Paper: We considered only peer-reviewed research papers.
-IC2: Primary Study: We included primary studies that directly addressed AI-OS integration.
-IC3: Relevance to AI and OS: Papers needed to discuss AI techniques enhancing OS or propose novel approaches for improving AI systems using OS methods.
-IC4: Publication Date: We restricted our selection to papers published within the past 5 years.
-EC1: Avoiding Duplicates: Only the most comprehensive or recent version of each study was included.
-EC2: Length Constraint: Papers shorter than 6 pages were excluded.
-EC3: Language: We considered papers written in English.
-As a result, 108 papers met our criteria and are within the scope of our research. Excluded papers merely mentioned the terms but did not delve into the actual intersection of OS and AI.
-Figure 1: Year Distribution of Paper Published
-3.3 Paper Analysis
-3.3.1 The main venues
-Figure 1 depicts the distribution of publication years for research 1.
-Notably, the number of papers published in the AI4OS and OS4AI domains indicates a growing trend, increased attention in this field.
-Notably, the number of papers published(including pre-print) in the field of integrating LLM and OS indicate a rapidly growing trend, 17 publications in this field since 2023, when powerful LLM emerged.
-Journals and conferences are listed in Tables 1 , the selection criterion is how many times they appear in our paper pool.
-International Conference on Architectural Support for Programming Languages and Operating Systems(ASPLOS) is the most prevalent conference.
-TABLE I: Conferences Distribution
-Conference Papers
-ASPLOS 22
-HPCA 11
-USENIX ATC 11
-OSDI 5
-SOSP 2
-HotOS 2
-APsys 2
-SOSP 2
-DAC 1
-CCS 1
-MobiCom 1
-ESEC/FSE 1
-ACM TURC 1
-ISSTA 1
-We first divide all the papers into AI4OS(38), OS4AI(53), LLM AS OS(9) and LLM4OS(8).
-4 RQ1: What OS sub-domains are researchers inclined to enhance using AI?
-Currently, OSes employ global, static policies based on heuristics.
-However,AI techniques, adaptable to different application behaviors, hold promise for outperforming these traditional policies.
-This research question explores and quantifies the diverse applications of AI approaches in enhancing or automating various OS tasks.
-4.1 AI for OS Tuning
-Growing OS complexity challenges configuration and decision-making, prompting the exploration of AI and ML for enhancements in auto-tuning tasks, including scheduling, energy efficiency, and memory management.
-Figure 2: General AI for OS Tuning Workflow
-Figure 2 showed the general workflow for AI tools to assist OS.
-Scheduling
-Scheduling, a core OS function, balances fairness, responsiveness, and throughput by efficiently managing resources.[2][16] Advancements through machine learning can refine ticket distribution, enable adaptability, and adjust scheduling variables precisely.
+3.1.2 Search database
+To gather all relevant articles related to our research, we must select suitable sources.
+For this purpose, we chose databases commonly used by researchers for conducting surveys and reviews.
+The selected sources are listed below:
+• ACM Digital Library (www.acm.org/dl)
+• IEEE Xplore (ieeexplore.ieee.org)
+• arXiv by Cornell University (www.arxiv.org)
+• Google Scholar (www.scholar.google.com)
+The search result from different data sources are listed in Table 1, and has a total of 824 papers.
+3.1.3 Inclusion Criteria and Exclusion Criteria
+Then following inclusion criteria (IC) and exclusion criteria (EC) are conducted:
+IC1: The work is publicated after 2019.
+IC2: Primary study.
+IC3: The work is about integrating AI to assist OS.
+EC1: The duplicate version of a selected paper (only the most comprehensive or recent version of each study should be included).
+EC2: The length of the paper is less than 6 pages.
+EC3: Paper written in a language other than English and Chinese.
+EC4: If an article is repeating in different resources, then the older duplicate was excluded.
+After applying these ICs and ECs, 88 research papers are selected. Detailed information on the selected papers is provided in B.
+[…]
+my root size/.style= minimum height=1cm, text width=1cm , my tnode size/.style= minimum height=2cm, text width=1cm
+{forest}
+[…]
+Figure 1: Literature Survey Tree
+4 Machine Learning Enhanced Operating Systems
+OSs manage a vast number of processes that continuously start and stop.
+These systems could benefit from analyzing historical execution behavior to optimize resource management, such as intelligently allocating resources based on usage patterns.
+However, OSs often rely on hard-coded heuristics that lack adaptability to changing usage scenarios (?).
+This highlights the need for mechanisms that can leverage the accumulated knowledge and patterns of user behavior to determine optimal configurations for operational timings, including thread scheduling, memory paging, and the frequency of buffer cache flushing.
+In this section, we will review the efforts documented in the literature that detail the application of ML methods to enhance OS functionalities.
+4.1 Scheduling
+Scheduling, a core mechanism for OSs, balances fairness, responsiveness, and throughput by efficiently managing resources (?, ?).
+The effectiveness of a scheduler directly impacts overall system performance, influencing factors such as task latency, resource utilization, and user experience.
+However, in the context of increasingly complex hardware architectures and diverse software workloads, traditional OS schedulers, which rely heavily on static, heuristic-based algorithms, often struggle to adapt to dynamic and heterogeneous environments.
+These limitations manifest as inefficiencies in handling workload variability, hardware resource contention, and energy optimization.
+To address these challenges, recent research has explored the integration of ML techniques into OS scheduling.
+By leveraging data-driven models, ML-based schedulers aim to automate decision-making processes and transcend the constraints of traditional approaches.
+These adaptive methods enhance scheduling flexibility, improve prediction accuracy for workload behaviors, and optimize resource allocation in real-time, thus addressing the evolving demands of modern computing systems.
+Nevertheless, the existing heuristic-based algorithms are difficult to achieve its most efficiency since they tend to focus on locally optimal solutions, often missing the global optimum due to limited exploration and static rules.
+These algorithms are typically tailored to specific scenarios and struggle to generalize to dynamic or complex scheduling problems with intricate dependencies.
+Their reliance on predefined rules and parameters can make them inflexible and sensitive to changes, leading to suboptimal performance in varying conditions.
+Moreover, heuristics lack scalability and fail to guarantee solution quality, making them unsuitable for applications requiring precision or adaptability.
+As a result, modern approaches like machine learning or metaheuristics are often preferred for more robust and adaptive scheduling solutions.
+In contrast to conventional heuristic-based approaches, generative models can also be utilized in scheduling to address challenges such as modeling dynamic and complex environments.
+Their ability to capture underlying patterns and properties from observed data makes them well-suited for simulating and predicting system behaviors.
+For instance, generative model-based approaches can effectively represent intricate workload patterns or resource contention scenarios, enabling more accurate modeling and decision-making.
+Additionally, these models excel in handling the complexity of modern computational environments, offering improved adaptability and precision compared to traditional methods.
+For OS in individual hosts, advancements through ML can refine ticket distribution, enable adaptability, and adjust scheduling variables precisely.
+Their purpose is to ensure fair allocation of time and memory among all processes while maintaining optimal system performance and enabling processes to complete their tasks as quickly as possible.
+Many researchers have proposed innovative work to optimise existing process scheduling algorithms, such as Completely Fair Scheduler (CFS).
+CFS ensures the fair distribution of CPU time among processes in the Linux kernel.
+However, vanilla CFS faces limitations in multicore environments, where its complexity increases significantly (?).
+Chen et al.
+(?) find that vanilla CFS approach maximizes the utilization of processing time but overlooks the contention for lower-level hardware resources.
+They present a ML-based resource-aware load balancer for the Linux kernel with a low-overhead method for collecting training data.
+Specifically, they present an ML model based on a multi-layer perceptron (MLP) that imitates the Linux’s CFS load balancer based on the collected training data and an in-kernel implementation of inference on the model.
+Using eBPF for dynamic tracing, an MLP model replicates CFS decisions, with in-kernel inference for real-time balancing, the model achieves high accuracy and small latency increase, demonstrating effective, low-overhead load management.
+This scheduler can be extended to bahve with higher efficiency.
+(?) port this algorithm and model to CUDA and place it in a kernel module to reduce its inference cost.
+Goodarzy et al.
+(?) also question vanilla CFS in its ability for proper allocation of CPU, memory, I/O, and network bandwidth.
+In response to this, the authors propose SmartOS, an operating system that automatically learns what tasks the user deems to be most important at that time.
+Based on the learned user preferences, SmartOS adjusts the allocation of system resources such as CPU, memory, I/O, and network bandwidth.
+It prioritizes the resources for the tasks that the user is currently focused on.
+The authors demonstrate an implementation of such a learning-based OS in Linux and present evaluation results showing that a reinforcement learning-based approach can rapidly learn and adjust system resources to meet user demands.
+(?) questions vanilla CFS in its struggles to effectively manage the excessive memory traffic generated by requests for DL-based services.
+The authors propose a memory-aware fair-share scheduling algorithm designed to reduce the susceptibility of QoS applications to memory-related interference.
+Besides process scheduling in the operating systems of individual hosts, ML techniques have been utilized to enhance resource allocation inside computer clusters (e.g., supercomputers).
 Springborg et al.
-[86] introduce Chronus, a Python application that collaborates with the Simple Linux Utility for Resource Management (SLURM) scheduler, prevalent in global supercomputers.
+(?) introduce Chronus, a Python application that collaborates with the Simple Linux Utility for Resource Management (SLURM) scheduler, prevalent in global supercomputers.
 Chronus executes comprehensive benchmark tests on HPC clusters, varying parameters like core numbers, processor speeds, and hyperthreading.
 Throughout, it logs performance metrics and energy use for each setup.
-This data trains a machine learning model to recognize patterns, aiming to forecast the most energy-efficient configuration for specific jobs and systems.
+This data trains a ML model to recognize patterns, aiming to forecast the most energy-efficient configuration for specific jobs and systems.
 Proven through notable energy savings in benchmarks, this approach holds potential for broader application in HPC systems, promoting eco-friendliness and financial prudence without sacrificing computational power.
-Chen et al.[3] presents a ML-based resource-aware load balancer for the Linux kernel with a low-overhead method for collecting training data,an ML model based on a multi-layer perceptron that imitates the Linux’s Completely Fair Scheduler (CFS) load balancer based on the collected training data and an in-kernel implementation of inference on the model.
-The authors argue that CFS approach maximizes the utilization of processing time but overlooks the contention for lower-level hardware resources.Using eBPF for dynamic tracing, an MLP model replicates CFS decisions, with in-kernel inference for real-time balancing, the model achieves high accuracy and small latency increase, demonstrating effective, low-overhead load management.
-Goodarzy et al.[14] also questioned CFS in its ability for proper allocation of CPU, memory, I/O, and network bandwidth.
-In response to this, the authors propose SmartOS, an operating system that automatically learns what tasks the user deems to be most important at that time.
-ased on the learned user preferences, SmartOS adjusts the allocation of system resources such as CPU, memory, I/O, and network bandwidth.
-It prioritizes the resources for the tasks that the user is currently focused on.
-The authors demonstrate an implementation of such a learning-based OS in Linux and present evaluation results showing that a reinforcement learning-based approach can rapidly learn and adjust system resources to meet user demands.
-Storage
-Storage systems, along with their associated OS components, are engineered to cater to a broad spectrum of applications and fluctuating workloads.
+From the aforementionend,Compared with traditional schedulers, new scheduling strategies $S$ are often developed because $J(S)$ evolves rapidly and is subject to numerous diverse constraints arising from the complex runtime context.
+The application of ML to scheduling represents a paradigm shift, enabling more intelligent and dynamic resource allocation as shown in Figure.2.
+By addressing limitations in traditional algorithms like CFS, ML-based solutions ensure fair resource distribution while optimizing system performance.
+These advancements, whether in individual hosts or large-scale HPC clusters, demonstrate the potential of ML to transform OS schedulers, balancing fairness, efficiency, and adaptability.
+Figure 2: Transition from simple scheduler to ML enhanced scheduler, due to complex running context
+Insights and Challenges:
+The integration of ML techniques into OS scheduling is driving a paradigm shift, enabling dynamic and intelligent resource allocation that exceeds the limitations of traditional heuristic-based approaches.
+ML-based schedulers demonstrate remarkable adaptability, capable of optimizing system performance in complex runtime environments by addressing challenges such as workload variability, resource contention, and energy efficiency.
+Unlike static schedulers, which struggle in current complex environments, these advanced techniques leverage data-driven models to improve prediction accuracy and decision-making.
+However, these advancements come with significant challenges that must be addressed for widespread adoption.
+The computational overhead of integrating ML models into real-time scheduling environments poses a critical limitation, as in-kernel inference must balance accuracy with minimal latency to ensure system performance is not compromised.
+Furthermore, reliance on extensive and high-quality training data for modeling workload patterns and resource utilization introduces complexity in dynamic and heterogeneous environments.
+Ensuring scalability across large-scale systems and generalizability to unseen scenarios remains difficult, particularly in multicore systems.
+Trade-offs between competing goals, such as fairness, throughput, latency, and energy efficiency, further complicate optimization, while embedding ML models into existing OS kernels or cluster management systems increases design and implementation complexity.
+Despite these challenges, the transformative potential of ML-based scheduling is clear, promising more intelligent, adaptable, and efficient resource management solutions for modern computing systems.
+4.2 Storage systems
+Storage systems, along with their associated OS components, are designed to handle a broad spectrum of applications and fluctuating workloads (?).
 The storage elements within the OS incorporate a range of heuristic algorithms, ensuring optimal performance and adaptability across diverse workloads.
-Predictable latency can be very useful for data-center systems serving interactive applications such as messaging and search.Cortez et al.[6] introduce LinnOS, a novel approach to managing SSD performance by incorporating a lightweight neural network within the operating system’s core.
+However, these approaches are often limited in their ability to capture the complex, nonlinear relationships between input parameters and performance outcomes.
+Predictability is a critical attribute of storage systems, as it ensures consistent and reliable performance across varying workloads and operational conditions.
+In modern computing environments, where storage systems must support a wide range of applications, from latency-sensitive tasks to high-throughput workloads, predictable behavior is essential for maintaining Quality of Service (QoS) guarantees and meeting Service Level Agreements (SLAs) (?, ?).
+With the increasing adoption of Solid-State Drives (SSDs) due to their low cost, high performance, and superior efficiency compared to traditional storage devices, SSDs have become a widely popular choice in storage systems (?).
+As a result, significant research efforts have been directed toward improving the predictability of SSD performance, aiming to address challenges and ensure reliable operation in diverse application scenarios.
+Cortez et al.
+(?) introduce LinnOS, a novel approach to managing SSD performance by incorporating a lightweight neural network within the operating system’s core.
 This neural network enables real-time inference of SSD performance at an extremely fine-grained level, specifically per-I/O operation, without requiring modifications to the underlying hardware or changes to file systems and applications.
-LinnOS profiles the latency of a large number of I/O operations submitted to the SSD, leveraging this data to train the neural network.
-By adopting a ”black-box” perspective towards the device, LinnOS learns and infers the SSD’s behavior to increase predictability, thereby empowering applications to anticipate whether their performance expectations will be met.
+By adopting a ’black-box’ perspective towards the device, LinnOS learns and infers the SSD’s behavior to increase predictability, empowering applications to anticipate whether their performance expectations will be met.
+Compared to traditional methods, LinnOS reduces I/O latency by up to 40% and improves throughput by up to 3× under high-concurrency workloads.
+Another critical aspect of SSD performance is the efficiency of Flash Translation Layers (FTLs), which manage the mapping between logical and physical addresses.
+Wang et al.
+(?) propose LearnedFTL, a machine learning-enhanced page-level FTL design that tackles address translation inefficiencies.
+This method models the non-uniform yet largely linear relationships between logical and physical page numbers, overcoming the irregularity in flash memory’s address space distribution.
+LearnedFTL integrates the training of learned indexes within the garbage collection (GC) process, continuously updating and refining the models.
+A bitmap prediction filter safeguards against potential model inaccuracies, ensuring correct mappings for address translation, thus minimizing double reads during random read accesses.
+Compared to state-of-the-art FTLs, LearnedFTL reduces address-translation-induced double reads by up to 55.5%, achieving a P99 tail latency improvement of 2.9× to 12.2× and an average throughput improvement of 5.5× to 8.2× under random workloads.
+Non-volatile memory (NVM) technologies face unique challenges, such as the write disturbance (WD) problem, which impacts reliability and endurance.
 Wu et al.
-[85] introduce LearnWD, a novel approach that harnesses the synergies of machine learning and out-of-place updates to effectively mitigate the write disturbance (WD) problem in NVM.
-At its core, LearnWD employs clustering algorithms to systematically categorize stale data based on their inherent error proneness.
-Upon receiving a write request, LearnWD meticulously assesses both the aggressiveness of incoming new data and the error vulnerability of the existing stale data.
-By thoughtfully leveraging this information, LearnWD strategically orchestrates write operations to proactively minimize the incidence of WD errors.
-Wang et al.
-[87] introduce LearnedFTL,a page-level Flash Translation Layer design that harnesses the power of ML techniques to significantly enhance the random read performance of flash-based SSDs.
-This method effectively models the non-uniform yet largely linear relationships between logical page numbers and physical page numbers, overcoming the inherent irregularity in flash memory’s address space distribution.LearnedFTL embeds the training of learned indexes within the GC process, allowing for the continuous updating and refinement of the learned models as the address mapping evolves.
-LearnedFTL also incorporates a bitmap prediction filter acting as a safeguard against potential model inaccuracies, validating the predicted PPNs and ensuring that only correct mappings are used for address translation.
-This innovative approach uniquely minimizes the occurrence of double reads resulting from address translation during random read accesses.
-Memory
-Traditional heuristic-based methods struggle due to the intricate hierarchy of memory levels, dynamic workloads, data-dependent behavior, and hardware constraints.
-These complexities make it difficult to optimize memory management effectively.
-Zhang et al.[17] address the challenges of address mapping in 3D-stacking memory, focusing on technologies like High-Bandwidth Memory and Hybrid Memory Cube.
-They propose Software-Defined Address Mapping, allowing user programs to directly control memory hardware.
-SDAM enables fine-grained data placement while leveraging chunk-based address mapping management.
-The system employs machine learning to identify access patterns, resulting in significant speedups compared to fixed address mapping systems.
-Lagar-Cavilla et al.[88] introduce a software-defined far memory system that has been successfully deployed across Google’s data centers since 2016.
-The authors propose a proactive software-defined approach that compresses cold memory pages to create a far memory tier in software.
-This approach uses the Linux kernel’s zswap mechanism to compress pages and store them in DRAM, effectively creating a far memory tier with low latency.
-The authors designed an autotuning system that employs a Gaussian Process Bandit to estimate the size of cold memory and the promotion rate under different configurations.
-This model is used to emulate the control algorithm offline and estimate the impact of different parameter settings.
-The authors conclude that their software-defined far memory system is effective in saving memory costs without impacting application performance.
-Rocha et al.
-[15] propose PredG, a Machine Learning framework designed to enhance the performance of graph processing.
-PredG aims to find the ideal thread and data mapping on Non-Uniform Memory Access systems.
-One of the key features of PredG is its ability to be agnostic to the input graph.
-It uses the available features of the graphs to train an Artificial Neural Network to perform predictions as new graphs arrive.
-This is done without any application execution after being trained, which makes it a powerful tool for optimizing graph execution on NUMA machines.
-This work is part of a broader trend in the research community towards automated graph machine learning, which seeks to discover the best hyper-parameter and neural architecture configuration for different graph tasks/data without manual design.
-4.2 AI for OS security
-With the widespread application of deep learning in recent years, using deep learning technologies for OS security has emerged, and the effectiveness of threat detection has been dramatically improved.
+(?) introduce LearnWD, a method that harnesses ML and out-of-place updates to mitigate the write disturbance (WD) problem in NVM.
+LearnWD employs clustering algorithms to categorize stale data based on their inherent error proneness.
+When a write request arrives, LearnWD analyzes the aggressor pattern of the new data and matches it to a stale block from the most suitable cluster, minimizing potential WD errors.
+By leveraging this speculative decision-making process, LearnWD reduces WD-prone writes while maintaining efficiency.
+Additionally, LearnWD requires only occasional retraining of its models to sustain its accuracy and performance.
+LearnWD is highly versatile and can be applied across various types of NVM, including single-level cell (SLC) and multi-level cell (MLC) configurations.
+Although the design is primarily optimized for phase-change memory (PCM), it is also extendable to other NVM technologies such as resistive RAM (ReRAM) and spin-transfer torque RAM (STT-RAM).
+Experimental results demonstrate that LearnWD achieves significant improvements: reducing WD errors by 20.1%, lowering write latency by 11.0%, and extending write endurance by 21.9% across 15 diverse real-world datasets.
+Figure 3: General ML for Storage Enhancement Workflow
+The aforementioned examples illustrate a typical pattern as shown in Figure.3, in which ML tools are often treated as a black box within storage systems to predict the future state of read/write requests.
+Once the decision-making process is complete, the operation that best aligns with the prediction is executed.
+Insights and Challenges:
+A significant trend in recent years is the integration of machine learning into storage systems, offering exciting opportunities to enhance performance and predictability in increasingly dynamic and complex environments.
+New type of storage hardware has stronger motivation for reserchers than those for schedulers.
+Unlike traditional heuristic-based methods, ML-driven solutions leverage data-driven models to capture non-linear relationships and adapt to diverse workloads in real-time.
+These approaches reduce manual intervention, improve resource utilization, and enable intelligent decision-making for tasks such as I/O scheduling, address translation, and error mitigation.
+Despite these advancements, several key challenges remain.
+First, the overhead associated with inference in ML models can limit their practicality, particularly in high-concurrency or resource-constrained storage systems.
+Storage systems has higher time requirements.
+Second, current methods often target specific hardware components or workload patterns, making it difficult to generalize their applicability across different storage technologies and use cases.
+Third, ensuring the robustness and reliability of ML models under unpredictable conditions, such as workload spikes or hardware failures, remains an open problem.
+4.3 Memory management
+Memory management and storage management are fundamental components of an operating system, each responsible for optimizing resource utilization and providing abstraction for different hardware resources.
+Memory management focuses on volatile primary memory (RAM), abstracting physical memory into virtual memory for fast, temporary storage.
+With the emergence of advanced hardware technologies such as 3D-stacked memory, significant improvements in memory bandwidth and latency have been realized.
+However, traditional heuristic-based memory address mapping algorithms often fail to fully leverage these advancements.
+To bridge this gap, Zhang et al.
+(?) propose Software-Defined Address Mapping (SDAM), a hardware-software co-design mechanism that dynamically adjusts address mapping strategies based on program access patterns, thereby maximizing hardware utilization in 3D memory.
+SDAM further integrates machine learning techniques to automatically identify and cluster access patterns of key program variables, enabling efficient and adaptive address mapping while minimizing overhead.
+This approach not only alleviates the burden on programmers but also allows the system to dynamically optimize memory performance for diverse workloads.
+In parallel with the advancements in high-performance memory technologies, new memory paradigms such as far memory are reshaping how memory capacity is scaled in modern systems (?).
+Addressing similar challenges of dynamic workloads and cost efficiency, Lagar-Cavilla et al.
+(?) introduce a software-defined far memory system for warehouse-scale computers (WSCs) to address increasing memory demands and cost challenges.
+By leveraging zswap, a Linux kernel mechanism, the system proactively compresses and migrates cold memory pages to create a flexible, low-cost far memory tier in software.
+It employs machine learning-based autotuning to optimize performance under dynamic workloads while meeting strict service-level objectives (SLOs).
+In the online section, the kernel manages memory with two key modules: kstaled and kreclaimd.
+kstaled tracks page ages by scanning page table entries, while kreclaimd reclaims cold pages exceeding the Cold Age Threshold, moving them to zswap to free DRAM.
+The Node Agent bridges the kernel and offline components by collecting statistics, adjusting the Cold Age Threshold dynamically, and exporting data for further analysis.
+In the offline section, an ML-based Autotuner uses far memory data and a Gaussian Process Bandit algorithm to optimize parameter configurations for memory savings and performance.
+The tuned parameters are fed back to the Node Agent, enabling continuous system optimization.
+Successfully deployed at Google since 2016, the system achieves significant memory cost savings (4–5% TCO reduction) with minimal performance impact, demonstrating practicality and scalability for large-scale production environments.
+Building on the theme of leveraging machine learning to optimize memory systems and workload performance, Rocha et al.
+(?) introduce PredG, a machine learning framework specifically designed to improve graph processing efficiency on Non-Uniform Memory Access (NUMA) systems.
+Unlike previous examples that focus on general-purpose optimization, PredG targets a specific application domain—graph processing.
+It identifies the optimal thread and data mappings for NUMA machines by utilizing graph features to train an artificial neural network for precise predictions.
+Notably, this framework eliminates the need for application execution after training, making it a powerful and efficient tool for optimizing graph workloads on NUMA architectures.
+This work aligns with the broader trend of automated graph machine learning, which aims to discover optimal hyper-parameters and neural architectures for diverse graph tasks and datasets without manual intervention.
+New deployment paradigms are also emerging in memory management, reflecting the evolving landscape of modern computing systems.
+Virtualization, a critical component of contemporary computing environments, introduces unique challenges and opportunities for optimizing memory and resource management.
+Continuing the exploration of applying machine learning techniques to these areas, Miura et al.
+(?) tackle the issue of cache management in virtualized environments, which involve both guest and host operating systems.
+In such setups, the second cache (managed by the host OS) often experiences poor locality of reference, making traditional algorithms such as LRU less effective.
+To address this, the authors propose leveraging LSTM, a recurrent neural network model designed for time-series data, to enhance cache replacement decisions in the second cache.
+Through simulations and real-world workload evaluations, their approach demonstrates significant improvements in cache hit ratios and I/O throughput.
+This work highlights the potential of machine learning to address inefficiencies in traditional memory and storage systems, showcasing how advanced models can adapt to complex, dynamic environments.
+It aligns with the broader trend of using ML to optimize memory management, offering innovative solutions for improving performance in virtualized systems.
+Insights and Challenges:
+A significant trend in recent years is the emergence of new hardware and deployment paradigms, which have opened up exciting opportunities for researchers to explore innovative methods.
+Unlike traditional heuristic-based approaches, software-defined memory management frameworks, such as SDAM and software-defined far memory systems—leverage machine learning to dynamically adapt to workload behaviors.
+This not only reduces manual intervention but also improves resource utilization.
+Despite significant advancement in memory management, several key challenges remain.
+First, many existing approaches incur high inference overhead, which can limit their practicality, particularly in resource-constrained or latency-sensitive environments.
+Second, current methods often focus on specific aspects of optimization and rely heavily on domain expertise for design and tuning.
+This lack of generality highlights the need for more versatile and intelligent solutions that can adapt to a wide range of hardware and workload configurations.
+Future research should prioritize developing low-overhead, general-purpose methods that balance performance optimization with reduced reliance on specialized knowledge, enabling more efficient and scalable memory management frameworks.
+4.4 Network
+The rapid development of the Internet and mobile communication industries has significantly increased the complexity of modern network systems.
+This evolution is characterized by more intricate infrastructures, a greater diversity of connected devices and resources, and increasingly dynamic network structures (?).
+To meet the demands of this ever-changing environment, network systems must become more intelligent and agile, adapting to diverse device requirements and unpredictable conditions.
+Artificial intelligence techniques, particularly deep neural networks (DNNs), have demonstrated significant potential in enhancing network traffic management.
+Kwon et al.
+(?) examine the deployment of DNNs for network traffic data classification, where DNNs are used to automatically classify real network traffic data collected from the ONOS (Open Network Operating System) platform.
+Experimental results with simple network topologies indicate that DNNs can be a promising approach to effective network packet classification.
+Additionally, it is confirmed that the deployment of DNNs for real network traffic data classification should consider both the data packets intended for delivery and those required for network maintenance, as the classification performance of DNNs significantly depends on the nature of the network traffic data.
+Another critical aspect of modern networking is ensuring optimal data transmission rates under varying conditions.
+Li et al.
+(?) propose a solution named Machine Learning Rate Adaptation (MLRA), which uses ML techniques to identify correlations between transmission rates, throughput performance, and link quality, thereby achieving high-efficiency adaptability for the IEEE 802.11ac standard.
+MLRA employs a two-level rate search procedure and a congestion detection mechanism to ensure the scalability and thoroughness of the algorithm.
+Experimental results show that, on commercial IEEE 802.11ac NICs, MLRA outperforms existing RA solutions with performance improvements ranging from approximately 133.0% to 658.0%, with minimal overhead.
+Beyond traffic and rate management, AI and Software-Defined Networking (SDN) are being combined to address broader challenges, such as network resilience during natural disasters.
+Veliyath et al.
+(?) introduce a novel approach to predicting, responding to, and rehabilitating from flood-based natural disasters using SDN, ML and image processing.
+It applies unsupervised KMeans clustering for optimal node distribution across satellite images and utilizes the supervised CatBoost algorithm with log transformation to forecast flood-affected areas.
+The study simulates an SDN controller using ONOS, known for its self-healing capabilities, next-generation device support, and high performance, and deploys a virtual network setup on ONOS using Mininet.
+Simulations show this approach can self-heal and enhance network performance.
+Insights and Challenges:
+The rapid evolution of mobile network systems, driven by increasingly complex infrastructure and dynamic network structures, demands intelligent and adaptive solutions for efficient management.
+The integration of AI and machine learning (ML) into network systems offers transformative potential.
+The researches highlights the potential of SDN with ML.
+SDN has ablitiy to flexibly adapt to different infrastructure and policies.
+Combining with ML tools, it demonstrate the versatility of AI in enhancing network resilience, efficiency and self-healing capabilities.
+These insights illustrate the growing role of AI-driven approaches in optimizing network performance, adaptability, and disaster recovery in increasingly dynamic and complex environments.
+Despite the promise of AI and ML in mobile network systems, several challenges remain.
+First, the performance of AI models, such as DNNs for traffic classification, heavily depends on the quality and diversity of the training data.
+The classification accuracy varies with the type of network traffic, requiring careful consideration of both delivery and maintenance packets.
+Second, while solutions like MLRA demonstrate significant performance gains, their scalability in highly dynamic and heterogeneous environments with diverse devices and traffic patterns is still a concern.
+Third, the computational overhead of deploying ML models, especially in real-time applications like traffic classification and disaster prediction, can strain resource-constrained network devices.
+Fourth, current mention methods still lack generalization ablitity across different platform and different targets.
+Addressing these challenges is crucial to fully realize the potential of AI in modern mobile network systems.
+4.5 Ops
+OS Ops are crucial for ensuring that systems function reliably and securely, thereby supporting the smooth operation of businesses and services that depend on them.
+Given the increasing complexity and scale of modern IT infrastructures, the potential for leveraging AI to improve OS Ops and maintenance is significant.
+Unlike the previously discussed sections, where ML modules are typically embedded within the OS to directly enhance scheduling, memory management, or storage systems, this section focuses on ML’s role as an external monitor.
+Here, ML operates outside the OS, analyzing system behaviors, logs, and workload patterns to make predictions and provide actionable insights.
+A common starting point for such monitoring efforts is the analysis of system logs, which are the most accessible resources for engineers seeking clues about an OS’s status.
+Building on this idea, Dusane et al.
+(?) introduce LogEA, a specialized tool for forensic investigations of Linux-based systems.
+LogEA uses natural language processing (NLP) and machine learning to analyze system logs, identifying security breaches, unauthorized access, and other anomalies.
+By detecting log messages with negative sentiment, LogEA flags and alerts users to potential events of interest, offering a targeted approach to anomaly detection.
+This focus on log analysis highlights the importance of leveraging existing OS outputs to uncover irregularities.
+Also based on logs, while LogEA focuses on forensic analysis and post-event investigation, Akram et al.
+(?) take a more proactive approach with Desh, a framework designed to diagnose and predict failures with short lead times.
+Desh applies a novel three-phase deep learning pipeline to logs, not only identifying chains of events that lead to failures but also predicting when and where failures are likely to occur.
+In the first phase, sequences of phrases leading to node failures are extracted, and LSTM Phase 1 is trained to identify and recognize such chains based on the training data.
+In the second phase, the chains formulated in Phase 1 are fed into LSTM Phase 2, which learns to account for the cumulative time differences ($\delta$ times) of the phrases within a failure chain, relative to the terminal phrase in that chain.
+During the inference phase, the chains learned in Phase 2 are used to estimate the lead times of future failures, along with predictions of the likely failure locations, based on test data that is completely independent of the training data.
+Desh utilizes an RNN with input/output layers and multiple hidden layers that form a stacked LSTM architecture, enabling effective training and prediction on system logs.
+Unlike LogEA, which alerts users to negative events retrospectively, Desh enables preemptive actions, such as migrating computation to healthy nodes, to minimize downtime.
+This shift from reactive forensics to proactive failure prediction showcases the versatility of ML in OS monitoring, as it can both analyze past events and anticipate future issues.
+Android is the most widespread smartphone operating system, and its popularity has attracted attackers who develop various malicious applications.
+On the defensive front, considerable research has been dedicated to identifying the types and states of Android applications based on their system-level behavior.
+Maasmi et al.
+(?) use a ML technique to detect whether an application is running in the foreground or not.
+This technique aims to enhance the accuracy of behavioral malware detection by providing informative priors or contextual metadata on the app’s state to malware identification models.
+The results indicate that a structured ML pipeline, which first identifies the app and then detects its mode, can achieve substantially higher accuracy compared to direct mode identification.
+In contrast to the log-based approaches above, Zhang et al.
+(?) propose a hardware-based solution for real-time workload forensics.
+Their method analyzes process execution flows by dividing them into consecutive frames and extracting Translation Lookaside Buffer (TLB) utilization features.
+These features are then fed into machine learning models to identify workloads at the granularity of individual processes.
+Unlike software-based workload forensics, which depend on OS services and can be compromised by software attacks, this hardware-based approach operates independently of the OS, offering a secure and efficient solution for monitoring workload behavior in real-time.
+This distinction highlights a critical difference: while tools like LogEA, Desh, and Maasmi’s pipeline leverage OS-level data for analysis, Zhang et al.’s method bypasses the OS entirely, focusing on hardware-level monitoring to enhance security and reliability.
+Insights and Challenges:
+Machine learning module as an external monitor for OS operations offers new opportunities to enhance reliability, security, and efficiency.
+Unlike embedded ML approaches, external monitoring uses logs, behaviors, and workload patterns to provide insights and predictions, enabling both forensic analysis and proactive failure prevention.
+However, challenges persist.
+Log-based methods depend on high-quality logs, and ensuring low overhead while maintaining real-time responsiveness can be difficult.
+Generalizability of ML models across different workloads and hardware configurations remains limited, requiring extensive retraining.
+Additionally, security and interpretability of ML solutions must be improved to prevent adversarial exploitation and ensure trust in critical systems.
+Monitoring is just one part of OS Ops.
+To truly reduce the burden on engineers, future AI-driven tools need to be more intelligent, general-purpose, and capable of self-healing, enabling systems to automatically detect, diagnose, and resolve issues without human intervention.
+4.6 Security
+OS security is one of the most prominent areas where ML tools are being employed to enhance functionality and robustness.
+ML-driven solutions are increasingly critical in addressing the evolving landscape of cyber threats, from malware detection to ransomware prevention and advanced persistent threats (APTs).
+These advancements leverage sophisticated algorithms to improve detection rates, reduce false positives, and provide proactive defense mechanisms across diverse platforms and attack types.
+APTs are among the most challenging cybersecurity threats, characterized by their sophistication and ability to remain undetected for extended periods.
+APTs often evade traditional anomaly detection methods by exploiting system vulnerabilities and employing deceptive techniques.
+Detecting APTs with ML is particularly difficult due to the rarity of relevant datasets and the significant imbalance in data, which complicates accurate detection.
+To address this, Benabderrahmane et al.
+(?) introduce AE-APT, a novel method utilizing AutoEncoder (AE) architectures ranging from basic to Transformer-based designs.
+By incorporating online learning-based anomaly detection, AE-APT enables real-time analysis and immediate responses to ongoing APTs, offering a promising solution to one of the most complex areas of cybersecurity.
+Malware detection is a key focus in AI-OS integration, particularly for Android, the most widely used mobile operating system.
 Qin et al.
-[35]presents MSNdroid, a novel malware detector designed specifically for Android applications, leverages a combination of native API calls, permissions, system API call features, and a Deep Belief Network.
-By applying deep learning techniques to native code features, MSNdroid effectively detects Android malware.
-This approach involves extracting features from a comprehensive dataset comprising malicious applications and benign applications.
-Notably, MSNDroid achieves an impressive accuracy while maintains an impressively low false-negative rate.
+(?) introduce MSNdroid, an advanced malware detector that leverages deep learning techniques to analyze native API calls, permissions, system API features, and other application-level data.
+By utilizing a Deep Belief Network and a comprehensive dataset of malicious and benign applications, MSNdroid achieves high detection accuracy with an impressively low false-negative rate, making it a reliable tool for combating Android malware.
+Similarly, Alzaylaee et al.
+(?) propose DL-Droid, a deep learning-based system that employs dynamic analysis to detect malicious Android applications.
+By generating inputs based on application states, DL-Droid achieves a detection rate of up to 97.8% using dynamic features alone and 99.6% when combining dynamic and static features.
+This approach demonstrates the power of combining feature types to improve detection performance while enhancing code coverage.
 De Wit et al.
-[39] emphasize the value of incorporating machine learning in malware detection strategies for Android platforms.
-By leveraging accessible hardware data and sophisticated classification techniques, the study demonstrates the feasibility of identifying malware with a reasonable degree of precision, highlighting the potential of app-specific metrics in enhancing detection rates.
-This research contributes to the growing body of knowledge on AI-integrated security measures within operating systems, particularly pertinent to the Android ecosystem.
-4.3 Findings
-In response to Research Question 1, which explores the inclination of researchers to utilize AI in enhancing specific OS sub-domains, our comprehensive analysis of existing literature highlights two prominent areas: auto-tuning and security.
-These domains have attracted considerable attention due to their potential for transformative improvements through AI integration.
-Auto-tuning emerges as a key area where AI is being employed to overcome the shortcomings of conventional heuristic-based methods.
-These traditional approaches often struggle to cope with the ever-evolving demands of modern computing environments.
-Researchers are harnessing machine learning to dynamically adjust OS parameters, such as scheduling, energy management, memory allocation, and storage optimization, aiming to boost system performance and resource efficiency.
-AI’s predictive abilities are particularly crucial in enhancing scheduling mechanisms to strike a balance between fairness, responsiveness, and throughput, effectively managing critical resources, ensuring they can handle the intricate and fluctuating workloads common in today’s computing landscapes.
-The security domain has witnessed a surge in the adoption of deep learning technologies to combat malware threats, showcasing substantial advancements in detection efficacy.
-AI-powered malware detectors have proven superior in terms of accuracy and reduced false negatives compared to traditional machine learning models.
-Furthermore, AI frameworks are being developed to create adversarial malware, underscoring AI’s dual role in fortifying OS security while also exposing vulnerabilities in AI-based detection systems.
-5 RQ2: How can OS be optimized to improve the efficiency of AI systems?
-AI accelerators are different from traditional hardware, affecting all aspects of system design, from data-center scale to single-chip scale.
-They also add high requirement for system architecture, management, and programming[13].
-Previous work has shown AI jobs critically demand high-speed I/O and low-latency and high-bandwidth data communication[18][21].Various attempt on hardware has been done to improve ML application performance,for exmaple,using newly NVMe SSD [89], relying on hardware FPGA for the I/O communication control instead of relying on OS-level interrupts that can significantly reduce both total I/O latency and its variance and algorithm level.
-Accessing hardware through the kernel introduces a performance bottleneck.
-To mitigate this bottleneck, one effective approach is to bypass the kernel altogether, enabling userspace programs to directly access hardware[20].
-Bateni et al.[89] presents NeuOS, a comprehensive system solution aimed at providing latency predictability for executing multi-DNN workloads in autonomous systems while simultaneously managing energy optimization and dynamic accuracy adjustments.
-It coordinates system- and application-level solutions intelligently to ensure that multiple DNN instances operate efficiently and meet their respective deadlines to guarantee latency predictability.
-It manages energy consumption by dynamically adjusting parameters such as voltage and frequency scaling to minimize energy usage without compromising the latency predictability or accuracy requirements.
-Based on specific system constraints, NeuOS adjusts the accuracy level of DNN computations in real-time.
-This allows for trade-offs between computational precision and resource efficiency, ensuring that the system operates within its given constraints while maintaining an acceptable level of performance.
-Wang et al.[21] showed what network for GPU AI remoting,a technique where the execution of GPU APIs is managed remotely through a network on a remote proxy instead of running GPU computations locally on the machine.
-The study takes a GPU-centric perspective to derive minimum latency and bandwidth requirements and aims to ensure that unmodified AI applications can run on remoting setups using commodity networking hardware without performance degradation.
-The paper introduces a novel theoretical framework that quantifies the minimum network requirements essential for efficient GPU API remoting.
-By formalizing the relationship between network latency, bandwidth, and remoting efficiency, this model provides foundational insights.
-Serizawa et al.
-[19]proposed an solution focused on I/O bandwidth.The method aims to improve the reading performance of large training datasets by using high-performance I/O storage devices.
-The authors discusses the problem of copying datasets between local storage and shared storage and proposes a solution to conceal the time spent on copying by overlapping the copying and reading of training data.
-5.1 Findings
-Addressing Research Question 2, which delves into the ways OSes can elevate the efficiency of AI systems, our analysis underscores the pivotal function of OSes in handling the distinctive needs of AI tasks.
-To optimize AI efficiency through the OS, several strategic approaches emerge as essential.
-Firstly, there is a need to tailor the OS to cater specifically to the requirements of AI accelerators.
-This includes facilitating high-speed I/O operations, minimizing latency in communications, and enabling autonomous operation of AI tasks.
-Secondly, the development of specialized runtime systems and schedulers becomes crucial, ensuring optimal allocation of resources and efficient execution of AI processes.
-Thirdly,optimizing I/O bandwidth further enhances the performance of AI applications.
-Collectively, these strategies form a comprehensive framework for improving the efficiency and effectiveness of AI systems through optimized OS integration.
-6 RQ3: How and which AI techniques are typically used to improve OS?
-To answer RQ3, we review all studies that investigate AI techniques applied in OS in this section.
-Among 38 AI4OS previous work,those did not clear clarify the AI tools excluded,we list the AI tools that are used more than twice.
-In general, RF, RNN (LSTM), KNN, RL, MLP,etc.
-are the most widely-used techniques.
-6.1 Tools analysis
+(?) further explore malware detection on Android by incorporating hardware-level data and classification techniques, showing the potential of app-specific metrics to boost detection rates and contribute to the growing field of AI-enhanced OS security.
+Ransomware remains one of the most critical cybersecurity threats, targeting systems worldwide and causing substantial financial and data losses.
+Wihar et al.
+(?) propose a novel approach to ransomware detection using an LSTM-based model capable of analyzing system call sequences in real-time.
+This model identifies anomalous behavior indicative of ransomware activities with high accuracy and low false-positive rates.
+Integrated into the Linux kernel, this method facilitates seamless communication and classification while maintaining efficient resource utilization.
+Experimental evaluations show that the LSTM model provides a proactive defense mechanism against ransomware threats, significantly enhancing the security framework of Linux systems through its adaptive learning capabilities.
+Extending beyond detection, Jurečková et al.
+(?) propose a machine learning-based approach for the online clustering of malicious samples into malware families.
+Using a weighted k-nearest neighbor classifier for known families and an online k-means algorithm for emerging malware, their method achieves cluster purities ranging from 90.20% to 93.34%.
+This approach focuses on the static analysis of portable executable files for the Windows operating system, offering valuable insights for categorizing malware and understanding evolving threats.
+Insights and Challenges:
+Machine learning is transforming OS security by offering advanced capabilities in threat detection, behavior analysis, and adaptive defense.
+Unlike traditional rule-based systems, ML excels at identifying complex patterns in large, dynamic datasets, enabling the detection of previously unknown or evolving threats.
+By analyzing system behaviors, network traffic, and user activity, ML can not only identify anomalies but also predict potential attacks, allowing for proactive mitigation.
+Furthermore, ML’s ability to integrate static and dynamic features provides deeper insights into system vulnerabilities and attack surfaces, enhancing overall security frameworks.
+However, the deployment of ML in OS security reveals critical insights: robust models require diverse and high-quality training data, as well as mechanisms to adapt to the ever-changing nature of cyber threats.
+Real-time learning and decision-making are increasingly essential, as static models struggle to keep pace with sophisticated attackers.
+Moreover, the integration of ML into OS workflows must balance accuracy, speed, and resource efficiency to be practical in real-world environments.
+These advancements suggest that ML has the potential to shift OS security from reactive to proactive, but achieving this requires addressing challenges related to scalability, data scarcity, and adversarial robustness.
+4.7 Analysis
+4.7.1 Typical Pattern:
+Figure 4 showed the general workflow for ML tools to assist OS.
+Figure 4: General ML for OS Enhancement Workflow
+The workflow for enhancing operating systems through machine learning is structured into two distinct phases: training and deployment.
+In the training phase, diverse datasets are collected from key system components, including the CPU, memory, storage, and I/O operations, serving as the foundation for model development.
+This raw data undergoes extensive preprocessing and transformation to ensure it is properly structured and optimized for training purposes.
+Advanced machine learning techniques are then employed to develop specialized models designed to address specific tasks, leveraging methods such as XGBoost, and various regression techniques.
+In the deployment phase, the trained models are integrated into the operating system and applied to real-time operational data.
+By analyzing this incoming data, the system uses the predictive capabilities of the models to optimize performance dynamically.
+This phase involves fine-tuning configurations and generating actionable insights to improve key areas, including memory segment optimization, task prioritization, and dynamic resource reconfiguration.
+These enhancements illustrate the transformative potential of machine learning when systematically applied to operating systems, enabling more efficient, adaptive, and intelligent resource management.
+4.7.2 How to choose proper algorithms to improve OS?
+Among the previous studies surveyed, we figure out some popular algorithms, include Random Forests (RF/DT), Recurrent Neural Networks (RNN/LSTM), K-Nearest Neighbors (KNN), Reinforcement Learning (RL) and Multi-Layer Perceptrons (MLP).
 RF(DT)
-A decision tree is a type of machine learning model used when the relationship between a set of predictor variables and a response variable is non-linear,while random forest is essentially a collection of decision trees.
+A decision tree is a type of ML model used when the relationship between a set of predictor variables and a response variable is non-linear,while random forest is essentially a collection of decision trees.
 It is quick to fit to a dataset and easy to interpret.
 Chowdhury et al.
-[32] used RF as one if the models to accurately detect the attack from the network traffic.
+(?) used RF as one if the models to accurately detect the attack from the network traffic.
 To construct the random forest classifier, the authors employed the Random Forest Regressor, which serves as both a regressor and a meta estimator.
 It accomplishes this by fitting multiple decision trees to different subsets of the dataset.
-For our specific model, we opted for a forest containing 1000 trees.Ahmed et al.
-[23] used RF to build a device fitness model,based on the Dataset collected during runtime and statically.
-De Wit et al.[39] trained a statistical classifier able to recognize malware signatures in any log data collected on a smartphone.
-The classifier was trained, cross-validated, and tested using the dataset described above and RF classifier had a better performance.Metzge el at.[49] used RF to get a optimal kernel runtime switching slice size.
+For this specific model, the authors opted for a forest containing 1000 trees.
+Ahmed et al.
+(?) used RF to build a device fitness model,based on the Dataset collected during runtime and statically.
+De Wit et al.
+(?) trained a statistical classifier able to recognize malware signatures in any log data collected on a smartphone.
+The classifier was trained, cross-validated, and tested using the dataset described above and RF classifier had a better performance.
+Metzge el at.
+(?) used RF to get a optimal kernel runtime switching slice size.
 The model is a random forest regressor with 50 decision trees with a depth of two for the GPU model.
-Ongun et al.[36] used RF to get the probability of a command being malicious,based on labels dataset.
+Ongun et al.
+(?) used RF to get the probability of a command being malicious,based on labels dataset.
+Sun et al.
+(?) used RF to filter storage I/O parameters demonstrates that tuning only the important parameters can nearly achieve the high performance obtained by tuning all parameters.
+Known for their versatility in handling non-linear relationships and minimal tuning requirements, RF/DT are used for tasks such as attack detection and configuration settings.
+They are chosen for their ability to quickly fit to datasets and provide interpretable models.
 RNN(LSTM)
@@
 They have feedback connections, allowing them to maintain an internal state or memory.
-Each step in an RNN processes an input and updates its hidden state based on the current input and the previous hidden state.While LSTM ,a type of RNN architecture with a more complex cell structure,were introduced to address the vanishing gradient problem.
-Motivated by the problem that exploiting 3D-stacking memory’s performance is challenging because bandwidth utilization heavily depends on address mapping in the memory controller, Zhang et al.[17] used a software-defined address mapping, allowing user programs to directly control low-level memory hardware in an intelligent and fine-grained manner.LSTM is used to in a method to get access pattern information to select an address mapping.
+Each step in an RNN processes an input and updates its hidden state based on the current input and the previous hidden state.
+While LSTM, a type of RNN architecture with a more complex cell structure,were introduced to address the vanishing gradient problem.
+Motivated by the problem that exploiting 3D-stacking memory’s performance is challenging because bandwidth utilization heavily depends on address mapping in the memory controller, Zhang et al.
+(?) used a software-defined address mapping, allowing user programs to directly control low-level memory hardware in an intelligent and fine-grained manner.
+LSTM is used to in a method to get access pattern information to select an address mapping.
 It identify the major variables that significantly contribute to external memory access and have a substantial impact on memory traffic and data movement.
-WordEmbedding
-Word embeddings provide a way to achieve this by mapping words to dense vectors in a multi-dimensional space.
-Fu et al.[34] proposed a AI-based approach to help under-resourced security analysts to find, detect,and localize vulnerabilities.
-They utilized a word-level Clang tokenizer with a copy mechanism.
-This tokenizer broke down a C function into a sequence of tokens.Then word embedding was used to generate vector representations for each token in the sequence, capturing the semantic information among the input tokens.
-And further classifier was , Ongun et al.[36] explore techniques for representing command-line text using word embedding methods.Based on this,they devise ensemble boosting classifiers to differentiate between malicious and non-malicious commands.
+Miura et al.
+(?) proposed to apply LSTM for time series data, for cache replacement in the second cache.
+Designed to handle sequential data, RNN/LSTM are particularly suited for tasks that require capturing temporal dependencies, such as managing memory access patterns in 3D-stacking memory systems.
 KNN
-KNN is a supervised machine learning algorithm used for both classification and regression tasks.
+KNN is a supervised ML algorithm used for both classification and regression tasks.
 The fundamental idea behind KNN is simple: neighbors influence each other.
@@
 It is widely applicable in pattern recognition, data mining, and intrusion detection.
-Yang et al[42] introduce a KNN-based machine learning algorithms can accurately predict the Turnaround-time(TaT) of a process.
+Yang et al.
+(?) introduce a KNN-based ML algorithms can accurately predict the Turnaround-time (TaT) of a process.
 It can effectively reduce the TaT of the process and reduce the number of process context switches.
@@
 An MLP is a type of feedforward neural network used for supervised learning tasks, such as classification and regression.
-Chen et al.[3] argues that traditional Linux CFS scheduler maximizes the utilization of processing time but overlooks the contention for lower-level hardware resources and try to solve the above problem using an ML-based resource-aware load balancer.They employed supervised imitation learning to replace a portion of its internal logic with an MLP model.
+Chen et al.
+(?) argues that traditional Linux CFS scheduler maximizes the utilization of processing time but overlooks the contention for lower-level hardware resources and try to solve the above problem using an ML-based resource-aware load balancer.They employed supervised imitation learning to replace a portion of its internal logic with an MLP model.
 This trained MLP model emulates the kernel’s load balancing decisions.
 MLP is chosen because this current work doesn’t require a very complex model and MLP has a relatively simple implementation compared to the other models.
-Based on this work,Qiu et al.[78] propose the concept of reconfigurable kernel datapaths that enables kernels to self-optimize dynamically to reduce the cost of kernel.
+Based on this work,Qiu et al.
+(?) propose the concept of reconfigurable kernel datapaths that enables kernels to self-optimize dynamically to reduce the cost of kernel.
 The authors also used MLP ML model that can mimic Linux CFS decisions.
-6.2 Findings
-In addressing Research Question 3, which explores the methods by which AI techniques are utilized to enhance OS, our analysis reveals a variety of approaches tailored to meet distinct OS requirements.
-These AI techniques are strategically chosen to address a spectrum of OS challenges, encompassing security enhancements, performance forecasting, resource management, and process scheduling.
-The widespread adoption of these methods attests to their versatility and underscores the profound influence of AI in augmenting OS functionality and performance.
-This integration of AI into core OS components not only boosts operational efficiency but also paves the way for more intelligent and adaptable systems, capable of meeting the evolving demands of modern computing environments.
-7 RQ4: How and what novel OS architectures create opportunities for AI4OS?
+5 Large Language Model Enhanced Operating Systems
+The rapid development of LLMs has introduced transformative capabilities in natural language understanding and generation, making them a promising tool for advancing operating system (OS) design and interaction.
+LLMs, such as GPT-4 (?), LLaMA (?), and Gemini (?), have demonstrated remarkable proficiency in processing and generating human-like text, understanding complex instructions, and adapting to a wide variety of tasks with minimal or no task-specific training.
+These models are built on advances in deep learning architectures, particularly Transformer-based designs, and trained on vast datasets that enable them to capture context, semantics, and nuanced relationships between concepts.
+One of the defining features of LLMs is their ability to act as general-purpose reasoning and interaction tools.
+Unlike traditional machine learning models, which are often task-specific, LLMs are highly adaptable, capable of handling a broad range of tasks through natural language interfaces.
+Their ability to understand intent, provide detailed explanations, and generate contextualized responses makes them an ideal complement to OS functionalities.
+As operating systems serve as the backbone for user-computer interaction, integrating LLMs can enable more intuitive, conversational interfaces, where users can communicate their needs and preferences in natural language rather than relying on rigid command structures or graphical user interfaces.
+The integration of LLMs into OSs presents a significant opportunity to enhance the user experience and overall system functionality (?, ?).
+LLMs can transform the way users interact with and manage their computing environments (?).
+5.1 Development
+LLMs possess powerful natural language processing capabilities, making them highly effective at understanding historical documentation on human experiences with operating systems.
+This ability is particularly valuable in the era of the Internet of Everything, where engineers invest significant effort in adapting hardware and applications to achieve optimal resource utilization and efficiency.
+Kamath et al.’s LLaMaS (?) and Chen et al.’s AutoOS (?) both utilize LLMs to address challenges in operating system configuration optimization, but they focus on different domains and employ distinct methodologies.
+LLaMaS addresses the complexities of adapting to diverse hardware environments by using LLMs to interpret plain text hardware specifications and recommend optimized OS strategies.
+Its architecture comprises a frontend and a backend: the frontend deciphers system characteristics and converts them into actionable embeddings, which the backend then uses for real-time OS decisions, such as managing memory tier data movement.
+In an experiment, LLaMaS demonstrated its capability by using ChatGPT to dynamically adjust memory allocation for CPU and GPU tasks based on workload demands.
+By leveraging LLMs’ zero-shot learning capabilities, LLaMaS eliminates the need for manual adjustments or device-specific coding, significantly reducing administrative and development costs.
+In contrast, AutoOS tackles the challenge of Linux kernel configuration optimization, a critical task for AIoT devices with constrained hardware resources.
+AutoOS models the kernel configuration space as a dynamic tree structure and employs a state-machine-based framework to traverse, prune, propose, and correct configuration options.
+The LLM frontend identifies performance-critical options, while the backend applies these refinements to boost system performance.
+Experiments showed that AutoOS could enhance UnixBench scores by up to 25% compared to vendor-provided defaults, achieving superior results with minimal human input.
+By automating the debugging and refinement of kernel configurations, AutoOS reduces the time and expertise required for manual optimization, addressing the complexity of configuring over 15,000 kernel options.
+Both LLaMaS and AutoOS leverage LLMs to simplify traditionally complex OS tasks, minimizing the need for manual tuning and expert knowledge.
+They also demonstrate the ability of LLMs to adapt dynamically to changing requirements.
+Moreover, both frameworks utilize LLMs’ reasoning and zero-shot learning capabilities to autonomously make decisions, showcasing their effectiveness in reducing development and administrative overhead.
+Another important and obvious area where LLMs excel is coding.
+As a type of textual data, code can be considered another form of human language, making LLMs particularly well-suited for tasks involving programming.
+This highlights the crucial role LLMs can play in improving software development processes, from code generation to debugging and vulnerability repair.
+For instance, SecRepair (?) is a system specifically designed to address the challenge of identifying and repairing code vulnerabilities.
+Powered by an LLM, it integrates reinforcement learning and a semantic reward mechanism to enhance its performance.
+To support its training, the authors have compiled and released a comprehensive instruction-based vulnerability dataset, which serves as a key resource for building a robust model for vulnerability analysis.
+Furthermore, SecRepair employs a reinforcement learning technique with semantic rewards to generate concise and contextually appropriate code commit comments.
+This approach, inspired by how humans resolve code issues, not only automates the repair process but also provides developers with a clear understanding of the vulnerabilities and their resolutions, ultimately streamlining software development and improving code security.
+Insights and Challenges:
+LLMs have the potential to significantly enhance software development processes, including tasks such as code generation, debugging, and vulnerability detection, by reducing manual effort and improving overall efficiency.
+The development and release of comprehensive, instruction-based datasets are essential for training models capable of addressing complex tasks like vulnerability analysis and repair.
+However, several challenges remain.
+For instance, mainstream operating systems like Linux consist of vast codebases with deeply coupled structures, making it difficult for current LLMs to fully capture all necessary modifications when proposing adaptations or optimizations.
+While LLMs can provide efficient suggestions, they may overlook related areas in the code that also require changes.
+Additionally, training a new LLM model demands significant time and computational resources, posing challenges in scenarios where engineers frequently switch to new hardware platforms.
+It is often impractical and uneconomical to train a fresh model for each new hardware adaptation, highlighting the need for more efficient and scalable solutions.
+5.2 User Interaction
+LLMs are transforming user interaction by enabling more natural, intelligent, and personalized experiences across various systems.
+As the critical bridge between hardware and software, the operating system plays a pivotal role in shaping the user interaction experience, which in turn defines how users engage with and perceive technology.
+Researchers are now striving to bring this revolutionary shift into the realm of operating systems, aiming to redefine how users interact with computing environments.
+Personalized services can meet specific user needs, enhancing satisfaction and loyalty.
+Furthermore, interactive systems can understand users’ perspectives and needs through well-designed interfaces and experiences.
+Zhang et al.
+(?) explores the potential applications of large language models, ML, and interaction design in recommendation systems and operating systems.
+By integrating these technologies, more intelligent and personalized services can be provided, meeting user needs and promoting continuous improvement and optimization of products, which holds great value for both recommendation research and user experience applications.
+Some researchers are exploring deeper structural reforms in OSs through the integration of LLMs.
+The AIOS-Agent ecosystem (?, ?) positions large language models (LLMs) as an ”operating system with a soul,” embedding them into the OS kernel to enable intelligent decision-making and resource management in Figure.6.
+The LLM’s context window functions as memory, handling relevant data, while external storage acts as a file system with advanced retrieval capabilities.
+Hardware and software are treated as peripherals and libraries, facilitating seamless agent-environment interaction.
+Natural language serves as the primary programming interface, democratizing software development.
+This ecosystem supports both single- and multi-agent applications capable of executing a wide range of tasks.
+Building on the AIOS concept, various studies have further enriched the ecosystem with additional components.
+Figure 5: Architecture for AIOS
+Figure 6: Architecture for LSFS
+Expanding on this foundation, the CoRE (Code Representation and Execution) system (?) introduces a unified framework for natural language programming, pseudo-code programming, and flow programming.
+CoRE leverages LLMs as interpreters to execute logically structured programs described in natural language, bridging the gap between human intent and machine execution.
+By defining a structured syntax, the CoRE language allows users to articulate program logic in a manner that is both accessible to non-programmers and precise enough for computational interpretation.
+Programs are decomposed into components such as step names, types, instructions, and connections, enabling clear and logical execution.
+Within the AIOS ecosystem, the LLM-based Semantic File System (LSFS) (?) in figure.6 redefines traditional file management by integrating LLMs for natural language-driven interactions and semantic understanding of file content.
+Unlike conventional systems that rely on metadata such as file names and timestamps, LSFS employs embedding vectors to index files based on their content, enabling advanced functionalities like semantic retrieval, grouping, and version control.
+Users can perform complex operations—such as file updates and change summarization—through simple natural language prompts, significantly streamlining file management.
+Key features include semantic indexing, natural language parsing for file operations, real-time change monitoring, and robust safety mechanisms to prevent unintended actions.
+Seamlessly integrated into the AIOS ecosystem, LSFS enhances intelligent multi-agent collaboration and supports capabilities such as task scheduling and knowledge integration.
+This innovation transforms file systems into intelligent, user-friendly components that are essential for modern AI-driven environments.
+Insights and Challenges:
+LLMs are driving a paradigm shift in user interaction by enabling natural, intuitive, and personalized experiences.
+Their integration into OSs redefines how users interact with hardware and software, making technology more accessible and user-centric.
+Users can interact with computing environments in a more intuitive manner, leading to increased satisfaction and loyalty.
+The shift towards personalized services allows systems to adapt to specific user needs, leveraging insights from interaction data to continuously refine and optimize user experiences.
+For instance, the work of researchers like Zhang et al.
+(?) highlights how LLMs can enhance recommendation systems.
+This trend underscores the importance of understanding users’ perspectives, enabling systems to respond dynamically to their needs through well-crafted interfaces.
+Moreover, innovative frameworks such as AIOS-Agent and CoRE exemplify the potential of LLMs in redefining OS functionality.
+By embedding LLMs into the OS kernel and employing natural language as a programming interface, these systems democratize software development and empower users—regardless of technical background—to engage with technology more effectively.
+Despite these promising developments, several challenges remain in the integration of LLMs within operating systems.
+One significant concern is the lack of interpretability associated with LLMs, which can impact their reliability.
+Users and developers may struggle to understand how LLMs arrive at specific outputs, leading to a lack of trust in their decisions and recommendations.
+This opacity can complicate troubleshooting and diminish user confidence, particularly in critical applications where understanding the rationale behind actions is essential.
+Additionally, there are challenges related to the scalability of these systems.
+As LLMs require significant computational resources, optimizing their performance in diverse hardware environments remains a critical hurdle.
+This is particularly important in ensuring that systems can deliver real-time responses without compromising efficiency.
+Moreover, while LLMs can enhance user interaction, they also raise questions about privacy and data security.
+The reliance on user data to personalize experiences necessitates robust safeguards to protect sensitive information and maintain user confidentiality.
+5.3 Analysis
+Existing research has explored various ways in which LLMs can be leveraged within OS architectures.
+LLaMaS, for instance, addresses the complexities of diverse computing environments by using LLMs to ease OS challenges in hardware adaptation and resource management.
+AIOS-Agent ecosystem treats LLMs as an ”operating system with soul”, embedding them in the OS kernel for intelligent decision-making and resource allocation.
+MemGPT introduces virtual context management to overcome the fixed-length context window limitation in LLMs, enabling them to access information beyond their immediate capacity.
+Additionally, systems like SecRepair and ChronoCTI demonstrate how LLMs can be employed to address challenges in software vulnerability repair and cybersecurity threat analysis, respectively.
+5.3.1 Differnce compared to ML-enhanced OS:
+LLMs and traditional ML techniques enhance OS in distinct ways, reflecting their unique capabilities and methodologies.
+Traditional ML techniques are often more task-specific and rely on structured data for training.
+These methods are typically employed for optimization tasks, such as system performance tuning, resource allocation, and configuration management.
+Additionally, ML models are frequently viewed as black boxes, making it challenging for users to understand the underlying principles and the relationship between the input training data and the resulting outputs.
+This lack of interpretability can hinder trust and usability, especially in critical applications where comprehending the rationale behind decisions is essential.
+In contrast, LLMs primarily focus on natural language understanding and generation, enabling more intuitive user interactions through conversational interfaces.
+They excel in interpreting user intent, managing complex tasks through natural language commands, and providing personalized experiences based on user input.
+This integration allows LLMs to transform how users engage with the OS, making technology more accessible and user-friendly.
+While there has been extensive research exploring the integration of traditional ML techniques into operating systems, where they effectively delve into the complexities of system performance, resource allocation, and configuration management, the role of LLMs remains more superficial.
+LLMs primarily function as an additional layer that enhances user interaction with the OS through natural language processing, enabling more intuitive and conversational interfaces.
+6 How do novel OS architectures create opportunities for intelligent OS?
 Comprehensive research has revealed that utilizing the standard kernel pathway for hardware interaction results in notable efficiency losses, rendering conventional approaches less than ideal for the current AI-driven environment.
@@
 This confluence of advancements signals a pivotal moment in the trajectory of computing, where the synergy between AI and OS architectures could redefine the boundaries of what is achievable in high-performance computing.
-7.1 Kernel-bypass OS Structure for AI
-Prior research has successfully demonstrated the potential of leveraging hardware acceleration for machine learning within the kernel space, showcasing the feasibility of such an approach[4].
+6.1 Kernel-bypass OS Structure
+Prior research has successfully demonstrated the potential of leveraging hardware acceleration for ML within the kernel space, showcasing the feasibility of such an approach (?).
 This study introducES an API remoting system, which facilitates access to specialized accelerator interfaces for kernel space applications.
@@
 Raza et al.
-[53] propose integrating unikernel optimizations into Linux, known for creating secure, compact OS images for single applications.
+(?) propose integrating unikernel optimizations into Linux, known for creating secure, compact OS images for single applications.
 Unikernel Linux (UKL) reduces the number of executed instructions and improves instructions-per-cycle efficiency.
@@
 It introduces a configuration option that allows a single, optimized process to link directly with the kernel, bypassing the traditional system call overhead, and significantly cuts latency for system call payloads, showcasing the benefits of unikernel-inspired enhancements in refining Linux’s performance.
-Cadden et al.[59] introduces Serverless Execution via Unikernel SnapShots (SEUSS), a method that leverages unikernel snapshots for rapid deployment and high-density caching of serverless functions.
+Cadden et al.
+(?) introduces Serverless Execution via Unikernel SnapShots (SEUSS), a method that leverages unikernel snapshots for rapid deployment and high-density caching of serverless functions.
 The authors describe the use of unikernel contexts, which consist of a high-level language interpreter configured to import and execute function code, providing isolation and security.
@@
 One significant obstacle that OSs face in effectively leveraging the potential of AI lies in the absence of a comprehensive, universally applicable strategy for adapting AI technologies to the wide array of heterogeneous devices in use.
-The Demikernel project[52] unveils an OS architecture optimized for datacenter systems with microsecond-scale requirements, emphasizing low-latency I/O.
+The Demikernel project (?) unveils an OS architecture optimized for datacenter systems with microsecond-scale requirements, emphasizing low-latency I/O.
 Adopting a LibOS strategy, it side-steps the traditional kernel in I/O paths, significantly enhancing performance.
@@
 Skiadopoulos et al.
-[60] present DBOS as a superior alternative to conventional cluster OS components, offering comparable functionality but enhanced analytics and reduced code complexity.
+(?) present DBOS as a superior alternative to conventional cluster OS components, offering comparable functionality but enhanced analytics and reduced code complexity.
 DBOS matches current systems in scheduling, file management, and inter-process communication, yet excels in analytics and simplifies code through database query-based OS services.
@@
 DBOS’s integrated DBMS approach is especially advantageous for ML, delivering a cohesive platform for efficient resource management and analytics in large-scale distributed environments, adeptly managing parallel computation and workload dynamics across various hardware.
-Shan et al.[61] developed LegoOS—a splitkernel architecture for hardware disaggregation, decentralizing OS functions for scalable, distributed management.
+Shan et al.
+(?) developed LegoOS—a splitkernel architecture for hardware disaggregation, decentralizing OS functions for scalable, distributed management.
 It fundamentally breaks down traditional OS functionalities into loosely-coupled monitors that each run on and manage a distinct hardware component.
@@
 The splitkernel model in LegoOS allows for independent scaling of compute and memory resources, essential for ML scenarios where large datasets and complex models necessitate specialized hardware accelerators.
-7.4 Findings
-Addressing Research Question 4, our analysis highlights the transformative role of novel OS architectures in advancing AI4OS.
-Pioneering designs like LibOS and unikernels have sparked renewed excitement among researchers, demonstrating their potential to outshine traditional OSes.
-These cutting-edge architectures enhance the symbiosis between software and hardware, reducing latency to boost AI application responsiveness and real-time processing capabilities.
-Their streamlined construction optimizes resource allocation, ensuring efficient management of computational assets.
-Beyond performance, these architectures establish a secure foundation for AI deployment.
-By simplifying the environment, they minimize security risks, vital for AI applications managing sensitive data.
-DBOS and LegoOS exemplify this by simplifying distributed AI setup and maintenance, freeing experts to focus on algorithmic innovation.
-They also adapt nimbly to the variable demands and hardware diversity of modern ML, enhancing deployment efficiency.
-In essence, these advanced OS architectures elevate performance and security, fostering an innovative landscape for AI4OS.
-By catering to AI application requirements, they promise to rewrite the rules of computing, blending AI and OS synergies for smarter, more efficient, and secure computing futures.
-8 RQ5: How LLMs create opportunities for OS?
-The integration of LLMs into OSes presents a significant opportunity to enhance the user experience and overall system functionality[67][68].
-LLMs, with their advanced natural language processing capabilities,[62][63][64] can transform the way users interact with and manage their computing environments.[76] We simply divide these works into two big categories: LLM AS OS and LLM4OS.
-8.1 LLM AS OS
-”LLM AS OS” refers to the integration of LLMs into the core of an OS, effectively serving as the ”brain” of the system,make OS capable of understanding and responding to natural language commands, thereby enabling more intuitive and flexible human-computer interaction[70][69].
-Kamath et al.’s LLaMaS[54] addresses the complexities of diverse computing environments by leveraging LLMs to ease OS challenges in hardware adaptation and resource management.
-It adapts to new devices by interpreting plain text specs, recommending optimized OS tactics.
-The LLM frontend deciphers system characteristics, converting them into actionable embeddings.
-The backend uses these for on-the-fly OS decisions, like memory tier data movement.
-An experiment with ChatGPT demonstrated its skill in adjusting memory allocation for CPU and GPU tasks based on usage.
-Aimed at reducing administrative and research costs, LLaMaS autonomously aligns with hardware changes via LLMs’ innate zero-shot learning, eliminating the need for manual adjustments or detailed device-specific coding.
-AIOS-Agent ecosystem[55][56], where LLMs act essentially as an ”operating system with soul”.LLMs are embedded in the OS kernel for intelligent decision-making and resource allocation.
-Its context window acts as memory, managing relevant data, while external storage serves as a file system with enhanced retrieval capabilities.
-Hardware and software are treated as peripherals and libraries, enabling agent-environment interaction.
-Natural language becomes the primary programming interface, democratizing software development.
-This ecosystem supports single and multi-agent applications for executing a broad range of tasks.
-MemGPT[57] presents a solution to overcome the fixed-length context window limitation in LLMs, a hurdle for tasks demanding deep analysis of lengthy dialogues or extensive texts.
-Drawing parallels with traditional OS hierarchical memory systems, MemGPT introduces virtual context management, echoing virtual memory principles.
-It enables LLMs to access information beyond immediate capacity, mirroring OS memory management through strategic ’paging’.
-This system reacts to events like user messages, system alerts, and timers, appending pertinent data to the primary context buffer before processing.
-Additionally, it supports function chaining for uninterrupted, sequential task execution, accommodating intricate operations and long-term planning within the LLM’s context limitations.
-8.2 LLM4OS
-SecRepair[58], designed to address the challenge of identifying and repairing code vulnerabilities in software development.
-This system is powered by a LLM and incorporates reinforcement learning and semantic rewards to enhance its capabilities.To support the training process and prepare a robust model for vulnerability analysis, the authors have compiled and released a comprehensive instruction-based vulnerability dataset.They also propose a reinforcement learning technique with a semantic reward mechanism to generate concise and appropriate code commit comments.
-This technique is inspired by how humans fix code issues and provides developers with a clear understanding of the vulnerability and its resolution.
-Rahman et al.
-[71] introduce ChronoCTI, an automated pipeline for mining temporal dynamics between attacker actions from text-based accounts of cyber incidents.
-The focus is on pinpointing repetitive action sequences, termed temporal attack patterns, crucial for preemptive defense strategies by security professionals against impending cyber threats.
-ChronoCTI is anchored by a curated dataset linking sentences to adversary maneuvers, temporal linkages across 94 attack narratives, and large language models refined on cybersecurity topics.
-The research delineates a structured approach, harnessing cutting-edge language models, NLP methods, and ML techniques to decipher the temporal structure within attack stories.
-8.3 Findings
-Addressing Research Question 5, our analysis highlights the incorporation of LLMs within operating systems heralds a transformative era, reshaping system dynamics and enhancing user engagement.
-They can enrich user interfaces by facilitating natural language processing for voice commands and text inputs, leading to more intuitive and personalized user experiences.
-In the realm of development, LLMs can automate code generation and optimization, streamlining coding processes and reducing errors.
-Additionally, they contribute to dynamic system management by analyzing usage patterns to optimize resource allocation, boosting system responsiveness.
-LLMs also play a pivotal role in intelligent troubleshooting, diagnosing system issues more accurately and recommending preventive measures to minimize downtime.
-Furthermore, they bolster OS security by detecting anomalies indicative of malicious activities and assisting in secure configuration management.
-Lastly, LLMs act as intelligent aids for developers, offering insights into function queries, suggesting best practices, and even generating documentation, thus accelerating the development process and enhancing code quality.
-Through continuous learning, LLMs ensure that OS performance and efficiency improve over time, adapting to user needs and operational contexts.
+6.4 Analysis
+All these works agree that, facing diverse computing environments, traditional OSs have encountered significant challenges.
+They reflect the increasing importance of AI in computing and the need for OS architectures that can effectively support AI workloads.
+They all strive to adapt to Heterogeneous Devices and Diverse Computing Needs.
+They focus on transforming low-level structures for efficiency and performance optimization in different aspects.
+Unikernels aim to enhance performance by reducing overhead and streamlining interactions between software and hardware.
+LibOS and split-kernels primarily focus on simplified integration and streamlined architectures, emphasizing scalability and flexibility.
+These works together pave a solid foundation for AI to refactor the structures of OSs, enabling better alignment with modern computing demands.
+Our analysis reveals several key innovations in OS architecture that offer new possibilities for the integration and application of AI:
+Kernel-Bypass OS Structures:
+Researchers have demonstrated the potential of leveraging hardware acceleration for machine learning within the kernel space.
+For instance, an API remoting system allows kernel space applications to directly access specialized accelerator interfaces, enhancing the performance of AI applications.
+Library OS:
+Addressing the challenges of datacenter systems requiring microsecond-scale responses, the Demikernel project introduces an OS architecture optimized for such demands.
+This LibOS strategy bypasses traditional kernels in I/O paths, significantly improving performance.
+LibOS components utilize hardware acceleration to offload critical tasks, minimizing latency and maximizing throughput.
+Other Structures:
+DBOS presents itself as a superior alternative to conventional cluster OS components, offering comparable functionality with enhanced analytics and reduced code complexity.
+It efficiently implements low-latency transactions and ensures high availability, making it especially advantageous for ML in large-scale distributed environments.
+LegoOS, developed as a split-kernel architecture for hardware disaggregation, decentralizes OS functions for scalable, distributed management, allowing independent scaling of compute and memory resources.
+These innovative OS architectures not only enhance performance and security but also provide better support for the needs of AI applications.
+They ensure efficient management through simplified environments and optimized resource allocation, providing a secure foundation for AI deployment.
+DBOS and LegoOS, by simplifying distributed AI setup and maintenance, allow experts to focus on algorithmic innovation.
+These architectures adapt nimbly to the variable demands and hardware diversity of modern ML, enhancing deployment efficiency.
+In summary, these advanced OS architectures play a transformative role in advancing AI-OS integration.
+They cater to the requirements of AI applications, promising to rewrite the rules of computing by blending AI and OS synergies for smarter, more efficient, and secure computing futures.
+7 How can we define the different stages of AI-assisted OSs?
+Figure 7: Three stages of AI-OS integration
+The evolution of AI-OS integration can be delineated into three progressive stages as shown in Figure 7, each marking significant advancements in the integration and utilization of AI within the foundational software that manages computer hardware resources and provides common services for computer programs.
+7.1 AI-Powered OS
+AI-powered Operating Systems build upon traditional architectures by integrating AI tools to enhance user experiences and optimize resource management efficiency.
+These enhancements include providing applications with common, system-level intelligent services such as understanding user needs (?), learning habits and preferences, and recommending applications.
+Additionally, these OSs more efficient and intelligent resource management services that are independent of hardware type and distribution, facilitating cross-device application execution, cross-architecture application migration, and dynamic hardware allocation based on operational data analysis.
+Furthermore, they support real-time fault detection and repair, dynamic monitoring of system operational states, and real-time adjustment of internal policies to ensure optimal operational performance.
+Lastly, they leverage AI in software engineering to automate or correct operating system code, thereby improving the efficiency and quality of OS development via automation and intelligence.
+Currently, most previous jobs focus on this stages.
+7.2 AI-Refactored OS
+The transition to an AI-Refactored Operating System (OS) involves a profound architectural transformation where intelligent mechanisms are deeply integrated into the OS framework (?, ?, ?).
+This advancement leads to a sophisticated level of software intelligence, fostering the development of intelligent, modularized OS structures.
+This stage paves the way for leveraging AI to conduct a sophisticated analysis of assets derived from open-source operating systems.
+AI algorithms formulate methodologies for constructing micro-libraries and facilitate the decoupling of existing systems, thereby maximizing the utilization of the current operating system ecosystem.
+In alignment with the generation requirements of the OS, AI identifies optimal combinations or configurations of these micro-libraries.
+Additionally, AI enables precise sensing of the operational state of the OS, which supports self-adaptive optimization during runtime, fostering the autonomous growth of the OS.
+Moreover, AI-Refactored OS facilitates the strategic planning of coordinated evolutionary pathways among multiple operating systems, promoting a collective process of self-evolution.
+Ultimately, this approach establishes an ecological paradigm grounded in shared capabilities, positioning itself competitively in the ongoing technological race.
+7.3 AI-Driven OS
+The AI-Driven OS signifies a radical departure from conventional frameworks by placing AI at the core of OS.
+It replaces the traditional operational logic with AI-driven logic and employs AI to manage and schedule various computing resources uniformly (?, ?, ?).
+This phase may see AI wholly or partially supplanting the traditional kernel, replacing heuristic-based decision-making systems with AI-driven operations, and supporting auto-generation and evolution of kernel modules under AI guidance.
+Additionally, it utilizes AI for comprehensive scheduling of diverse computing resources, achieving unified management and orchestration.
 8 Opportunities for Future Work
-9.1 Optimized OS for AI
-The escalating sophistication of AI workloads, characterized by high-velocity I/O, minimal latency, and bandwidth-intensive requirements, is catalyzing an urgent need for AI-specialized OS optimization.
+During our research,we find that the escalating sophistication of AI workloads, characterized by high-velocity I/O, minimal latency, and bandwidth-intensive requirements, is catalyzing an urgent need for AI-specialized OS optimization (?).
 As AI permeates diverse sectors, the imperative for a finely calibrated infrastructure escalates, poised to support the burgeoning demands of next-generation applications.
@@
 These technological marvels, while delivering unrivaled performance, necessitate software ecosystems meticulously attuned to their singular attributes.
-In this multifaceted hardware ecosystem [20][23], the need for an OS that can seamlessly manage and orchestrate AI tasks across a spectrum of devices is more pressing than ever.
+In this multifaceted hardware ecosystem (?, ?), the need for an OS that can seamlessly manage and orchestrate AI tasks across a spectrum of devices is more pressing than ever.
 AI has transcended beyond the realm of traditional neural networks, with innovative models leading the change in natural language processing and graph neural networks adeptly handling intricate relational data.
@@
 Juggling performance, energy conservation, and scalability constitutes a multifaceted challenge, demanding an OS equipped for dynamic adjustment to cater to the singular requisites of every workload.
-Portatble OS for AI
+Heterogeneous adaption for AI
 In industrial settings, IoT devices generate massive amounts of data.
 Deploying deep learning models directly on edge devices (such as sensors, gateways, or edge servers) in equally important as relying solely on cloud-based processing.
-In a certain sense, deploying on edge devices can be a more challenging scenario because the computational power and bandwidth, of edge devices are more constrained[77][12][1].
+In a certain sense, deploying on edge devices can be a more challenging scenario because the computational power and bandwidth, of edge devices are more constrained (?, ?, ?).
 Due to the diversity of edge devices,current AI platforms’ lack of portability across different edge platforms hinders widespread adoption ,and bypassing the kernel and directly accessing hardware while adapting to various hardware configurations can be extremely challenging.
-Therefore, an operating system that can accommodate multiple hardware types and is specifically tailored for AI applications is a potential solution[7][8].
-Fast deployment and standard API interface
-Simplified deployment means that even non-experts can utilize AI applications effectively.
-Developers, data scientists, and industrial engineers can quickly integrate the framework into their existing edge devices without extensive configuration or manual setup.
-And a standard API interface ensures that software components can communicate seamlessly with each other.
+Therefore, an operating system that can accommodate multiple hardware types and is specifically tailored for AI applications is a potential solution (?, ?).
 Minimized OS consumption
-Resource-constrained environments, such as edge devices or embedded systems, often operate with tight constraints on memory, processing power, and energy while in scenarios like cloud computing or data centers, minimizing resource consumption directly impacts operational expenses.
-As OS and applications share the resources, the OS consumption determines the actual operational performance of the algorithmic model.
+Resource-constrained environments, such as edge devices or embedded systems, often operate with tight constraints on memory, processing power, and energy while in scenarios like cloud computing or data centers, minimizing resource consumption directly impacts operational expenses.(?) As OS and applications share the resources, the OS consumption determines the actual operational performance of the algorithmic model.
 So there is a need to design an efficient OS that can achieve small memory and fast training and inference.
 There are several promising approaches to achieve this.
-Apart from bypassing the kernel, unikernel OSes are designed to be fast and lightweight[11][32].
+Apart from bypassing the kernel, unikernel OSs are designed to be fast and lightweight (?, ?, ?).
 Privacy and security
-The problem of privacy preservation in the context of machine learning is quite different from that in traditional data privacy protection, as machine learning can act as both friend and foe.
+The problem of privacy preservation in the context of AI is quite different from that in traditional data privacy protection.
 Increasingly, edge devices are equipped with AI applications.
-This trend comes with privacy risks as models can leak information about their training data [24].
+This trend comes with privacy risks as models can leak information about their training data (?).
 OS need to find ways to provide safeguards for ensuring the confidentiality and integrity of its data and programs.
-9.2 Intelligent OS
-9.2.1 Potential for Intelligent LibOS
+8.1 Library OS based Intelligent OS
+Inspired by previous requirement for OS enhancement for a intelligent OS, we believe Library os is a potential solution for it.
 Figure 8: Monolithic application to Microservice: ① Microservices enable independent, scalable deployments ② Microservices provide fine-grained security controls ③ Microservice enables fine-grained flexible upgrading and replacement of individual microservice ④ Microservice enables service intelligent orchestration ⑤ Microservice allows agile and independent development for each service
-Figure 4: Monolithic OS to LibOS:①LibOS spans across multiple hosts.②LibOS offers flexible, granular security and kernel-bypass capabilities.③Only LibOS supports granular updates to individual components.④LibOS enables orchestrating components intelligently for adaptable system contexts.⑤LibOS facilitates agile, independent development of single components.
-A recently emerging trend of Internet-based software systems is “resource adaptive,” i.e., software systems should be robust and intelligent enough to the changes of heterogeneous resources[9], both physical and logical, provided by their running environment[26].
+Figure 9: Monolithic OS to LibOS: ① LibOS spans across multiple hosts.
+② LibOS offers flexible, granular security and kernel-bypass capabilities.
+③ Only LibOS supports granular updates to individual components.
+④ LibOS enables orchestrating components intelligently for adaptable system contexts.
+⑤ LibOS facilitates agile, independent development of single components.
+A recently emerging trend of Internet-based software systems is ‘resource adaptive’, thus software systems should be robust and intelligent enough to the changes of heterogeneous resources (?), both physical and logical, provided by their running environment (?).
 The key principle of such an OS is based on resource disaggregation, resource provisioning as a service, and learning-based resource scheduling and allocation.
 Meanwhile, the OS should leverage advanced machine/deep learning techniques to derive configurations and policies and automatically learn to tune itself and schedule resources.
-One general and efficient Spark tuning framework is proposed and applied it to Tencent’s data platform[5].
+One general and efficient Spark tuning framework is proposed and applied it to Tencent’s data platform (?).
 LibOS have risen to prominence as highly attractive options in scenarios that prioritize agility and robust security, thanks to their inherent minimalist and modular architectures.
@@
 Microservices excels notably in granular security, isolated updates, and performance optimization through dynamic service management.
-AI technologies have proven adept at governing microservices, enabling intelligent service orchestration and customization, as highlighted in research [65] and [66], setting the stage for advanced system control and optimization[81].LibOS similarly exhibit gains in agility, composability, enhanced isolation, and facilitated deployment across heterogeneous hardware.
-LibOS presents a more auspicious ground for intelligent OS.
-There are at least 5 ways where LibOS will benefit to enable a Intelligent OS:
+AI technologies have proven adept at governing microservices, enabling intelligent service orchestration and customization, as highlighted in research (?) and (?), setting the stage for advanced system control and optimization (?).
+LibOS similarly exhibit gains in agility, composability, enhanced isolation, and facilitated deployment across heterogeneous hardware.
+Heterogeneous deployment and minimized OS consumption
+In the IoT epoch, a surge in varied endpoint devices underscores the imperative for AI integration.
+Yet, conventional monolithic OS kernels struggle, weighed down by excessive resource demands and challenges in hardware acclimatization, hindering their agility in today’s rapidly transforming milieu (?).
+In alignment, LibOS surface as a beacon of hope for versatile, cross-platform deployment amid the AI revolution.
+The Demikernel project (?) substantiates LibOS’s flexibility, enabling applications to operate fluidly across disparate devices sans customization.
+In LibOS’s schema, hardware interface components are architected as pluggable modules, amplifying interoperability and malleability.The intrinsic modularity and adaptability simplify AI assimilation into a spectrum of devices, nurturing a more dynamic and reactive IoT framework.
 Isolation and security
-FlexOS[82] presents a modular LibOS framework, comprised of discrete elements that can be segregated using a spectrum of hardware safeguards, complemented by adaptable data sharing protocols and software fortification methods.
+FlexOS (?) presents a modular LibOS framework, comprised of discrete elements that can be segregated using a spectrum of hardware safeguards, complemented by adaptable data sharing protocols and software fortification methods.
 FlexOS adopts micro-libraries as fundamental units for segmentation.
@@
 Drawing from the microservices paradigm’s strengths in isolation and precise access governance, we posit that LibOS offers a strategic avenue to uphold data confidentiality during processing and safeguard overall system integrity.
-In the AI domain, where security and privacy are paramount[37].
+In the AI domain, where security and privacy are paramount(?).
 LibOS’s robust isolation features make them ideally suited for secure operation in environments with concurrent tasks of varying privacy sensitivity.
 The innate simplicity and modularity of LibOS create an ecosystem that effectively quarantines critical workloads, bolstering confidence and aligning with rigorous privacy statutes.
-Heterogeneous deployment
-In the IoT epoch, a surge in varied endpoint devices underscores the imperative for AI integration.
-Yet, conventional monolithic OS kernels struggle, weighed down by excessive resource demands and challenges in hardware acclimatization, hindering their agility in today’s rapidly transforming milieu.
-In alignment, LibOS surface as a beacon of hope for versatile, cross-platform deployment amid the AI revolution.
-The Demikernel project[52] substantiates LibOS’s flexibility, enabling applications to operate fluidly across disparate devices sans customization.
-In LibOS’s schema, hardware interface components are architected as pluggable modules, amplifying interoperability and malleability.The intrinsic modularity and adaptability simplify AI assimilation into a spectrum of devices, nurturing a more dynamic and reactive IoT framework.
 Intelligent OS orchestration
@@
 This capability serves as an inspiration for the realization that similar traffic control principles can be harnessed within the realm of LibOS to achieve intelligent resource management.
-In the context of demanding tasks such as machine learning training that consume vast amounts of resources, as well as in resource-constrained edge devices, meticulous oversight over resource allocation is paramount.
+In the context of demanding tasks such as ML training that consume vast amounts of resources, as well as in resource-constrained edge devices, meticulous oversight over resource allocation is paramount.
 This empowers dynamic, autonomous resource allocation, optimizing performance and efficiency across varied workloads and settings, perfectly suiting complex, evolving AI applications and infrastructures.
@@
 Intelligent OS development
-Microservices architecture heralds a seismic shift in software development paradigms, dismantling centralized structures in favor of agile, decentralized teams, each specializing in a unique microservice[83].
+Microservices architecture heralds a seismic shift in software development paradigms, dismantling centralized structures in favor of agile, decentralized teams, each specializing in a unique microservice (?).
 This transition boosts efficiency and reliability, as focused teams oversee discrete services, revolutionizing software engineering approaches.
 Contrastingly, monolithic systems’ vast scale and tangled dependencies escalate code generation complexities.
-Automating code in large, interconnected systems demands meticulous integration within a dense module network[84].
+Automating code in large, interconnected systems demands meticulous integration within a dense module network (?).
 LibOS, with their compact, independent nature, simplify code generation, offering fertile ground for efficient automation and AI-driven development.
-9.2.2 Multi-agent LLM for OS Development
-Previous research has demonstrated the efficacy of employing AI in optimizing, safeguarding, and evaluating system software including operating systems[10][22][28].
-Our study reveals a notable dearth of efforts dedicated to leveraging LLMs in the context of operating systems.
-Nonetheless, we identify at least one distinct domains where LLMs can potentially revolutionize the functioning and capabilities of OSes as Multi-agent LLM for OS Development.
-Figure 5: One example for Multi-agent LLM for OS Development
+Figure 10: One example for Applying Multi-agent LLM to Automate OS Development
 Recent breakthroughs in LLM-based multi-agent systems, building on the success of single LLMs as unified decision-makers, have unlocked new horizons in complex problem-solving and environment simulation.
 This progress signals the potential for multi-agent LLMs to revolutionize OS development.
-Studies like [79] and [80] showcase multi-agent LLM models’ efficacy in software development, from analysis to debugging.
-In the realm of OS development, the integration of multi-agent systems fosters a specialized, collaborative, and flexible environment.
-A constellation of agents, each dedicated to niches such as kernel architecture, driver engineering, or cybersecurity, collectively enhances project productivity[72][73][74][75].
+Studies like (?) and (?) showcase multi-agent LLM models’ efficacy in software development, from analysis to debugging.
+In the domain of OS development, the integration of multi-agent systems as shown in Figure 10, facilitates a specialized, collaborative, and adaptable environment.
+A constellation of agents, each dedicated to specific niches such as kernel architecture, driver engineering, or cybersecurity, collectively contributes to enhanced project productivity(?, ?, ?, ?).
 These agents facilitate knowledge exchange and consensus-building akin to human collaboration.
@@
 Certain agents conduct simulations to assess OS performance under varied scenarios, supplying empirical insights to inform judicious decision-making, preserving the OS’s technological relevance and competitiveness.
-9.2.3 LLM AS OS
+8.2 LLM Based Intelligent OS
+Ideas such as those presented in (?) and (?) propose innovative frameworks for current LLM-based operating systems or runtimes.
+However, in line with our definition of an AI-driven OS, we believe that the generalization capabilities of LLMs position them to function as various sub-modules of operating systems—such as schedulers or filesystems—effectively replacing traditional OS components.
+This approach can be realized without necessitating changes to existing applications or hardware adaptations, allowing for a smoother transition and integration.
+By leveraging the adaptive and conversational nature of LLMs, these systems could enhance resource management, task scheduling, and data handling in a more intuitive and efficient manner, ultimately transforming the landscape of operating systems while maintaining compatibility with current software and hardware infrastructures.
 Future research in the domain of LLMs integrated as OS will focus on following areas.
@@
 This personalization fosters a more engaging and productive user interface, tailored to unique usage patterns.
-Enhanced Security Measures
-LLMs contribute to OS security by analyzing data for threats, enforcing policies, and managing access.
-They can predict security incidents, provide real-time code analysis, and offer personalized security guidance, fortifying the system against vulnerabilities.
+Instructions Interporation
+Traditional operating systems link to applications through system calls, and if large models can accurately translate and understand these instructions, they indeed have the potential to replace the system call layer or even the entire operating system.
+By leveraging natural language processing and contextual understanding, large models could directly interpret application demands and generate corresponding system-level operations, thereby simplifying the interaction between users and the system.
+Such a replacement could enhance the system’s flexibility and usability while lowering the technical requirements for developers and users.
 9 Conclusion
-The synthesis of AI and OS is reshaping the landscape of contemporary computing, as evidenced by our meticulous review of 108 primary studies.
-AI’s integration into OS frameworks has catalyzed significant improvements in key areas like memory oversight, task coordination, and security vigilance.
-By leveraging intelligent decision-making beyond traditional rule-bound paradigms, AI bolsters resource efficiency and system agility.
-In tandem, advanced OS designs are morphing to better serve AI applications, especially as data-intensive workloads surge across various computing platforms.
-The OS’s evolving role in adeptly handling specialized hardware and optimizing data flow underscores its importance in sustaining AI software performance.
-The integration of LLMs into OS interfaces, as highlighted in our analysis of 17 studies, marks a pivotal advancement in user experience.
-Conversational interfaces enable seamless human-computer interaction, simplifying device management and personalizing user assistance through natural language processing.
-Despite the optimistic outlook, integrating AI into OS presents critical challenges.
-Safeguarding against security breaches and privacy infringements while maintaining efficient AI operations necessitates rigorous design considerations.
-Moreover, the demand for streamlined AI algorithms that can operate within constrained environments remains a high-priority research area.
-Future research agendas should concentrate on the domain of intelligent operating systems, capitalizing on the transformative potential they offer.
-Central to this endeavor is the exploration of LibOS.
-With its modular structure, LibOS facilitates the seamless incorporation of AI enhancements, supports granular control, fosters innovation, and upholds stringent security standards.
-LibOS stands as a robust foundation for the evolution of intelligent computing.
-Thus, advancing LibOS technologies should be a focal point for researchers aiming to realize the full scope of intelligent OS capabilities and drive the computing industry toward greater adaptability, efficiency, and resilience.
-Acknowledgments
-This work was supported in part by the Major Program of National Natural Science Foundation of Zhejiang(LD24F020014), and in part by the Zhejiang Pioneer (Jianbing) Project (2024C01032), and in part by the Key R&D Program of Ningbo(2023Z235), and in part by the Ningbo Yongjiang Talent Programme(2023A-198-G).
+In summary, this review has comprehensively explored the current state of AI-OS integration, highlighting significant advancements and identifying areas for further investigation.
+Key findings include the effective utilization of AI techniques, such as ML and large language models, to improve OS functionalities like memory management, process scheduling, and intrusion detection.
+The review underscores the importance of secure and efficient AI integration within OS frameworks, emphasizing the need for robust validation methods and evaluation metrics.
+Additionally, the potential of novel OS architectures that leverage AI capabilities is outlined, paving the way for intelligent, modular, adaptable, and self-aware operating systems.
+Future research should focus on addressing the challenges associated with integrating AI into OSs, including the development of more efficient training strategies, the creation of larger and more diverse datasets, and the improvement of AI models’ interpretability and explainability.
+Exploring the ethical implications of AI in OSs is also crucial, ensuring that the benefits of these technologies are balanced against potential risks.
+The convergence of AI and OS represents a pivotal arena for innovation in computing technology, with the development of intelligent OSs critical in shaping the future of computing.
+Acknowledgment
+This work was supported in part by the National Science Foundation of China under Grants (62472375,62125206), and in part by the Major Program of National Natural Science Foundation of Zhejiang (LD24F020014), and in part by the Zhejiang Pioneer (Jianbing) Project (2024C01032), and in part by the Key R&D Program of Ningbo (2023Z235), and in part by the Ningbo Yongjiang Talent Programme (2023A-198-G), and in part by the fund of Laboratory for Advanced Computing and Intelligence Engineering.
+A Tables
+Resource name Search string in Initial selection
+ACM Digital Library Abstract 210
+IEEE Xplore Digital Library Abstract 279
+arXiv by Cornell University Abstract 158
+Google Scholar Title 177
+Table 1: Search results from different data sources.
+Abbreviation Description
+MLP Multi-Layer Perceptron
+RNN Recurrent Neural Network
+CNN Convolutional Neural Network
+LSTM Long Short-Term Memory
+DT Decision Tree
+RF Random Forest
+GA Genetic algorithm
+LLM Large Language Model
+DNN Deep Neural Networks
+SDN Software Defined Network
+LRU least recently used
+Table 2: Abbreviations
+B Detailed information on selected papers
+Figure 11 depicts the distribution of publication years for research.
+The figure indicates that from 2019 to 2021, the number of published works was relatively low and stable.
+However, starting from the year 2023, there has been a steady increase in the number of publications.
+This could reflect an increasing level of activity in this field of research, as well as potential trends and growing interest in the subject.
+Figure 12(a) illustrates the distribution of publication venues for research.
+Notably, the ACM International Conference on Architectural Support for Programming Languages and Operating Systems (ASPLOS) has accepted a significantly greater number of publications.
+The distribution of countries is shown in Figure 12(b), where the USA, China, and India are the top three contributors to the total number of publications.
+Figure 11: Year Distribution of Paper Published
+(a) Venue Distribution of Paper Published
+(b) Country Distribution of Paper Published
+Declaration of generative AI and AI-assisted technologies in the writing process
+During the preparation of this work the author(s) used Qianwen and ChatGPT in order to improve the readability in the writing process.
+After using this tool/service, the author(s) reviewed and edited the content as needed and take(s) full responsibility for the content of the publication.